### 23 Nov 2023 v1

# Assumption-lean and Data-adaptive Post-Prediction Inference

Abstract
A primary challenge facing modern scientific research is the limited availability of gold-standard data which can be both costly and labor-intensive to obtain. With the rapid development of machine learning (ML), scientists have relied on ML algorithms to predict these gold-standard outcomes with easily obtained covariates. However, these predicted outcomes are often used directly in subsequent statistical analyses, ignoring imprecision and heterogeneity introduced by the prediction procedure. This will likely result in false positive findings and invalid scientific conclusions. In this work, we introduce an assumption-lean and data-adaptive Post-Prediction Inference (POP-Inf) procedure that allows valid and powerful inference based on ML-predicted outcomes. Its “assumption-lean” property guarantees reliable statistical inference without assumptions on the ML-prediction, for a wide range of statistical quantities. Its “data-adaptive” feature guarantees an efficiency gain over existing post-prediction inference methods, regardless of the accuracy of ML-prediction. We demonstrate the superiority and applicability of our method through simulations and large-scale genomic data.
1 Introduction
A fundamental challenge in modern scientific research is the acquisition of gold standard data (Wang et al.,, 2023). These data, with their high accuracy and reliability, are essential to the validity of scientific discoveries, but obtaining them is often costly and labor-intensive. Fortunately, the advent and rapid development of machine learning (ML) has made it possible to predict outcomes using accessible covariates (LeCun et al.,, 2015; He et al.,, 2016). A prominent example is AlphaFold (Jumper et al.,, 2021), which uses readily available protein amino acid sequences to accurately predict protein structures that traditionally require extensive experimental efforts to determine. This ML-based approach has demonstrated its potential to substantially reduce the time and resources required to measure gold standard data (Stokes et al.,, 2020; Cheng et al.,, 2023).
Despite these benefits, replacing gold standard data with ML-prediction introduces new challenges, particularly in maintaining the validity of downstream statistical analyses. The indiscriminate use of such predictions, without acknowledging their distinction from observed gold-standard data, can lead to biased results and misleading scientific conclusions (Wang et al.,, 2020). This issue is exemplified by the statistical analysis using imputed gene expression in the Genotype-Tissue Expression (GTEx) project.
Example 1 (Multi-tissue Prediction of Gene Expression in GTEx Project).
Integrating gene expression across tissues and cell types is critical to understanding the coordinated biological mechanisms that drive disease and characterize homeostasis (Lonsdale et al.,, 2013). The GTEx project is a comprehensive public resource that studies gene expression and regulation in multiple human tissues (GTEx Consortium et al.,, 2015). It aims to improve our understanding of gene regulation and thus provide insights into the biological mechanisms behind various diseases and traits. However, many types of tissue are difficult to access and are not collected from each GTEx individual. The percentage of individuals with missing gene expression ranges from 5% to 90% (median 47%) across all tissues in GTEx (GTEx Consortium,, 2020). This leads to an incomplete picture of an individual’s physiological state and requires effective methods for multi-tissue prediction. To address this problem, several approaches have been proposed to impute gene expression in the uncollected tissue (Wang et al.,, 2016; Viñas et al.,, 2023; Basu et al.,, 2021). However, after prediction, imputed gene expression has been treated directly as observed gene expression and used in subsequent statistical analysis for scientific discovery (Viñas et al.,, 2023), such as understanding the genetic basis of gene regulation. As demonstrated in (Wang et al.,, 2020), direct use of the estimation results is likely to produce false positives, leading to invalid scientific conclusions.
The problem of making inferences after acquiring ML predictions is not limited to the specific field in Example 1, but is instead a pervasive problem in various domains where ML is applied (Bullock et al.,, 2020). The inherent complexity and “black box” nature of many ML models adds layers of difficulty in validating the statistical inferences derived after prediction. This highlights the need for methods that can effectively bridge the gap between ML-prediction and the demands of scientifc rigor, ensuring that the integration of ML into research does not compromise the validity of scientific conclusions.
To address this problem, we introduce a simple, valid, powerful, and widely applicable approach called Post-Prediction Inference (POP-Inf) that is assumption-lean and data-adaptive. POP-Inf is designed to integrate ML-prediction with the observed gold standard data for a general estimation problem. We highlight two key features of our POP-Inf method:
  • Assumption-lean: The POP-Inf estimator is consistent and asymptotically normal, without making any assumptions about the ML prediction algorithm. This means that POP-Inf can be used with arbitrarily misspecified ML-prediction.
  • Data-adaptive: The POP-Inf estimator is adaptive to the accuracy of the ML-prediction. It utilizes more information from a “good” ML-prediction to reduce variance, and avoids variance inflation otherwise. It degenerates to many existing estimators under certain cases of the ML-prediction.
Our contributions are threefold: (i) We present a simple and data-adaptive method of post-prediction inference for an estimation problem defined through estimating equations, without any assumptions on the ML predictions or any parameterization of the true data generating process. (ii) We establish the consistent and asymptotic normality of our proposed estimator. We further demonstrate the optimality of our estimator in terms of asymptotic variance over a class of estimators that includes many existing estimators in the post-prediction inference. (iii) We connect our method with the theory of semiparametric efficiency. We derive the semiparametric efficiency bound and demonstrate the potential of our proposed method in achieving that bound. In addition, once the nuisance function cannot be correctly specified, an efficient influence function-guided estimator with estimated nuisances plugged in can be less efficient than the proposed POP-Inf procedure.
The rest of this paper is organized as follows. In Section 2, we formulate our problem and illustrate our method with mean estimation to build an intuition. Next in Section 3, we introduce a general protocol for applying POP-Inf to estimands defined through estimating equations. Further, we establish the asymptotic properties of the POP-Inf estimator and describe the estimation procedure. We build the connection with the semiparametric efficiency theory in Section 4. Through extensive simulations (Section 5) and real data analysis (Section 6), we validate our theoretical claims and demonstrate the practical utility and effectiveness of the POP-Inf estimator. This paper concludes with a discussion in Section 7.
2 Problem Formulation
Let $Y$ be a scalar outcome variable and ${\bf X}$ be a $d$-dimensional covariate variable. The scientific interest is to estimate a $q$-dimensional parameter ${\bm{\theta}}$ defined through an estimating equation,
$\mathbb{E}\{\bm{\psi}(Y,{\bf X};{\bm{\theta}})\}={\bf 0},$ (1)
where $\bm{\psi}(\cdot,\cdot;{\bm{\theta}})$ is a user-defined function. Such a definition of ${\bm{\theta}}$ is very general including outcome mean, outcome quantile, least squares coefficients, or any other specific quantities of interest involving both ${\bf x}$ and $y$, including the unique minimizer of a loss function and the maximizer of a criterion function (Van der Vaart,, 2000) ; see Section 2.1 for an example of estimating the outcome mean.
We observe a random sample where only a small subset is labeled with outcome $Y$. In addition, we observe auxiliary variable ${\bf Z}$ that is predictive of the outcome $Y$. Note that ${\bf Z}$ could be a subset of ${\bf X}$ and vice versa; see Remark 1. Denote that sample as ${\cal L}\cup{\cal U}$ where =${\cal L}=\{(y_{i},{\bf x}_{i},{\bf z}_{i}),i=1,\cdots,n\}$ and ${\cal U}=\{({\bf x}_{i},{\bf z}_{i}),i={n+1},\cdots,N+n\}$. Assume that ${\cal L}$ and ${\cal U}$ are independent and their marginal distributions of $({\bf X},{\bf Z})$ are the same. Let $n/N\to\rho$ as $n\to\infty$ and $N\to\infty$ and let $\rho=\infty$ if there is no unlabeled data.
We consider the availability of an external and independent prediction algorithm $\widehat{f}(\cdot)$ on ${\bf z}$ that produces predictions $\widehat{f}({\bf z})$ for the outcome variable $Y$. We assume that the operating characteristics of $\widehat{f}$ are unknown to the end-user, and the data used to fit $\widehat{f}$ are unavailable. Thus $\widehat{f}$ is considered a black box function. In addition, this $\widehat{f}$ is can be incorrect and even biased for predicting $y$.
With access to this black-box $\widehat{f}$, our primary goal is to improve efficiency while making inference on the parameter ${\bm{\theta}}$. We refer to our problem as post-prediction inference in that the inference is made after predictions having been made independently (Wang et al.,, 2020). Unlike existing work that provides tools to estimate the true “$f$” or assuming it can be well approximated, we make no assumptions on $\widehat{f}$. The resulting protocol can benefit researches where training a prediction algorithm is unrealistic (e.g., training ChatGPT), and provide guidelines for scientists who are not equipped with computational resources or knowledge about building prediction algorithms.
Remark 1.
(The consideration of ${\bf X}$ and ${\bf Z}$)
We deliberately split the covariate into two (possibly overlapping) parts, ${\bf Z}$ and ${\bf X}$, where the former is predictive of the outcome $Y$ and the latter appears in the estimating equation (1). They could be the same, but a wide range of applications suggest they are different. When ${\bf Z}$ is not a subset of ${\bf X}$, the information it carries about the outcome $Y$ may not be informative for the estimation problem (1). Hence, a better prediction algorithm $\widehat{f}$ does not necessarily correspond to a better estimate of ${\bm{\theta}}$. As justified above, this work does not attempt to improve the prediction algorithm, but to empower the inference on the parameter of interest, given any arbitrary black-box $\widehat{f}$.
In the followings, Section 2.1 illustrates the key idea of the proposed method by an example of estimating outcome mean. Section 2.2 reviews related work. After that, in Section 3, we introduce the proposed framework for the estimation problem defined in equation (1).
2.1 Example: estimating the outcome mean
We provide the intuition of our approach using mean estimation. Consider $U(y,{\bf x};{\bm{\theta}})=y-\theta$. The resulting estimand is $\theta=\mathbb{E}(Y)$, the outcome mean. For illustration propose, we consider ${\bf z}={\bf x}$. Then the classic approach calculates an estimator from the sample average of observed outcomes,
$\widehat{\theta}_{\textnormal{C}}=\dfrac{1}{n}\sum_{i=1}^{n}y_{i}.$ (2)
This estimator is unbiased, but can have high variance due to the small size of labeled data. To increase efficiency with the ML-predicted outcomes in unlabeled data, we propose to introduce another unbiased estimator of zero with a weight scalar $\omega$,
$\widehat{\theta}_{\textnormal{PopInf}}(\omega)=\dfrac{1}{n}\sum_{i=1}^{n}y_{i}+\omega\left[\dfrac{1}{N}\sum_{i=n+1}^{n+N}\widehat{f}({\bf z}_{i})-\dfrac{1}{n}\sum_{i=1}^{n}\widehat{f}({\bf z}_{i})\right]$ (3)
We refer to this estimator as the POP-Inf estimator.
Both the classic estimator and recently-introduced prediction-powered (PP) estimator (Angelopoulos et al., 2023a, ) can be viewed as special cases of the POP-Inf. The classic estimator is the POP-Inf estimator with $\omega=0$. The PP estimator is the POP-Inf estimator with $\omega=1$:
$\widehat{\theta}_{\textnormal{PP}}=\dfrac{1}{n}\sum_{i=1}^{n}y_{i}+\left[\dfrac{1}{N}\sum_{i=n+1}^{n+N}\widehat{f}({\bf z}_{i})-\dfrac{1}{n}\sum_{i=1}^{n}\widehat{f}({\bf z}_{i})\right].$ (4)
As detailed in Section 4, another candidate estimator is guided from the efficient influence function (EIF). We refer to this estimator as $*{}^{*}$-based where $*{}^{*}$ represents the fact that the nuisance function is plugged in as an estimate,
$\widehat{\theta}_{\textnormal{EIF}^{*}}=\dfrac{1}{n}\sum_{i=1}^{n}y_{i}+\dfrac{N}{N+n}\left[\dfrac{1}{N}\sum_{i=n+1}^{n+N}\widehat{f}({\bf z}_{i})-\dfrac{1}{n}\sum_{i=1}^{n}\widehat{f}({\bf z}_{i})\right].$ (5)
It is also a special case of the POP-Inf estimator with $\omega=N/(n+N)$. This estimator achieves the semiparametric efficiency bound when $\widehat{f}$ approximates its true counterpart at a sufficiently fast rate.
Since all these estimators are summations of an unbiased estimator of the population mean and an unbiased estimator of zero, they are consistent for the population mean. In addition, Wald-type confidence intervals can be constructed using the following asymptotic variance,
$\underbrace{\frac{\textnormal{Var}_{l}(Y)}{n}+\underbrace{\left(\omega^{2}\frac{\textnormal{Var}_{l}[\widehat{f}]}{n}-2\omega\frac{\textnormal{Cov}_{l}[Y,\widehat{f}]}{n}+\omega^{2}\frac{\textnormal{Var}_{u}[\widehat{f}]}{N}\right)}_{\text{Additional terms}}}_{\text{{POP-Inf}\leavevmode\nobreak\ interval}},$ (6)
where $\textnormal{Var}_{l}[\cdot]$ and $\textnormal{Cov}_{l}[\cdot,\cdot]$ represent the variance and covariance on the labeled, respectively, and $\textnormal{Var}_{u}[\cdot]$ is the variance on the unlabeled. Note that although $\textnormal{Var}_{l}[\cdot]$ and $\textnormal{Var}_{u}[\cdot]$ are essentially the same, we still use different notations to indicate their different sample analogs that will guide the practice.
To guarantee that the asymptotic variance (6) is no larger than the classic approach regardless of the quality of $\widehat{f}$, the additional terms need to be no larger than zero. A key insight is that these additional terms are a quadratic function of $\omega$,
$q(\omega)=\omega^{2}\left\{\frac{\textnormal{Var}_{l}[\widehat{f}]}{n}+\frac{\textnormal{Var}_{u}[\widehat{f}]}{N}\right\}-2\omega\frac{\textnormal{Cov}_{l}[Y,\widehat{f}]}{n},$ (7)
which achieves its minimum at $\omega_{0}=\textnormal{Cov}_{l}[Y,\widehat{f}({\bf z})]/\{\textnormal{Var}_{l}[\widehat{f}]+n\textnormal{Var}_{l}[\widehat{f}]/N\}$ with minimum value
$q(\omega_{0})=\dfrac{-\textnormal{Cov}_{l}[Y,\widehat{f}]}{n\textnormal{Var}_{l}[\widehat{f}]+n^{2}\textnormal{Var}_{u}[\widehat{f}]/N}\leq 0,$ (8)
where the equality holds if and only if $\textnormal{Cov}_{l}[Y,\widehat{f}]=0$. In this case, our POP-Inf estimator is no less efficient than the classic estimator. In contrast, both PP and $*{}^{*}$-based estimator may be less efficient than the classic estimator, when the ML predictions are not accurate. As we shall see in Section 3, similar patterns appear in estimating the general parameter defined in (1). In addition, the additional terms are below zero if and only if $\textnormal{Cov}_{l}[Y,\widehat{f}]>\omega(\textnormal{Var}_{l}[\widehat{f}]+n\textnormal{Var}_{u}[\widehat{f}]/N)/2$. This indicates that for efficiency improvement, the prediction accuracy of the ML algorithm cannot be low.
Remark 2 (Data-Adaptive Feature).
We provide an intuition for the “data-adaptive” feature of POP-Inf. When the ML algorithm is purely random, $\textnormal{Cov}_{l}[Y,\widehat{f}]=0$ and hence the optimal weight $\omega_{0}=0$. Therefore, POP-Inf degenerates to the classic estimator using labeled data only. Such choice suits the expectation that one should not use this ML-prediction. On the other hand, if the ML-prediction is perfect in that the predictions are exactly the true outcomes, $\textnormal{Cov}_{l}[Y,\widehat{f}]=\textnormal{Var}_{l}[\widehat{f}]=\textnormal{Var}_{l}[Y]$ and hence the optimal weight is $\omega_{0}=N/(N+n)$. In this case, POP-Inf degenerates to the EIF-based estimator that achieves the semiparametric efficiency bound. This is also intuitive since perfect ML-prediction can be treated as gold-standard data and the estimator should utilize all the (measured and predicted) information and thus is efficient.
Figure 1 displays numeric results for estimating outcome mean and ordinary least squares coefficients with details contained in Section 5. With a common $\widehat{f}$ obtained from an independent sample, we construct the POP-Inf estimator, the $*{}^{*}$-based estimator, the PP estimator, and the classic estimator. On the x-axis, each $\gamma$ corresponds to a data generating system which roughly speaking quantifies the capability of $\widehat{f}$ for predicting $y$: the higher $\gamma$ is, the better $\widehat{f}$ would be. On the y-axis, left panels display the coverage rates; right panels coordinate the ratio of averaged confidence interval widths across $1000$ simulations between the classic estimator and one of the competing estimators.
All methods yield satisfying coverage rates centered around the nominal level $95\%$, but their efficiency can be quite different. From right panels, the $*{}^{*}$-based estimator and the PP estimator perform comparably, both less efficient than the classic approach when $\gamma$ is small, and more efficient as $\gamma$ becomes close to $0.5$. On the other hand, the POP-Inf estimator is always equally or more efficient than the classic approach despite the value of $\gamma$. In addition, the POP-Inf estimator is more efficient than the $*{}^{*}$-based and PP estimators across $\gamma\in[0,0.5]$. Additional results in Section 5 corroborate such patterns. From empirical evidence, the POP-Inf estimator is desirable in practice.
Figure 1: Relative efficiency and coverage rates in simulations with $\gamma\in[0,0.5]$, $n=250$ and $N=20n$. In each panel, the x-axis coordinates the data generating parameter $\gamma$ detailed in Section 5. Left panels plot the ratios between averaged standard errors of a competing method and the classic approach, with dashed horizontal line intersects the y-axis at $1$. A ratio lower than one represents a smaller standard error than the classic approach. Right panels plot coverage rates with dashed horizontal line intersects the y-axis at the nominal level $0.95$.
2.2 Related work
Our setting is closely related to a recent method called prediction-powered (PP) inference (Angelopoulos et al., 2023a, ), which elicits valid inference with gold standard labels and arbitrary ML predictions. It corresponds to a special case of our POP-Inf estimator where $\omega=1$ (or a vector of all one’s). Both theoretical and numeric comparisons suggest that the PP estimator is less efficient than the proposed POP-Inf estimator.
Our proposal of boosting an estimator with an unbiased estimator of zero has been a long-standing idea. One famous estimator is the augmented inverse probability weighting estimator in the literature of missing data (Robins et al.,, 1994) and causal inference (Robins,, 2000), in which a consistent inverse probability weighting estimator is augmented by a weighted residual term with mean zero, resulting in efficiency gain (Hahn,, 1998). In fact, it reaches the semiparametric efficiency bound (Bickel et al.,, 1993; Hahn,, 1998) when nuisance functions are correctly specified or approximated sufficiently fast. Although the pursuit of efficiency under correct specifications differs from our main focus, we establish the efficiency bound of our estimation problem (1) in Section 4 and illustrate the connection between the POP-Inf estimator and the efficient influence function in an example in Section 2.1.
The idea of using unlabeled data fits into the broader context of improving efficiency with auxiliary data, which has been widely applied in survey estimation (Breidt and Opsomer,, 2017), missing data (Robins et al.,, 1994), measurement error models (Chen et al.,, 2005), causal inference with surrogate outcomes (Kallus and Mao,, 2020), and semi-supervised learning (Wang and Shen,, 2007; Chakrabortty and Cai,, 2018). The literature of semi-supervised learning has adopted a similar data structure with our problem in which labeled data are accompanied with unlabeled data (possibly with surrogate outcomes). Recently, Chakrabortty and Cai, (2018) have considered efficiency improvement on linear models under possible model misspecifications. After that, estimation methods have been proposed for mean estimation (Zhang et al.,, 2019), best linear predictor estimation (Azriel et al.,, 2022), and the general M-estimation problems (Song et al.,, 2023). Our work is distinct from this line of research in that we consider post-prediction inference with any “black-box” ML predictions. We aim at efficiency improvement with no assumptions on $\widehat{f}$ and for generally defined parameters.
3 POP-Inf
In this section, we introduce an assumption-lean and data-adaptive post-prediction inference procedure that guarantees variance reductions for a general estimand defined in (1). We study asymptotic properties of the proposed POP-Inf estimator in Section 3.1 and provide an estimation algorithm in Section 3.2. Examples on regression models are provided in Section 3.3.
Bearing in mind that $\widehat{f}$ could be incorrect, one of the most straightforward approaches to estimating ${\bm{\theta}}$ is simply ignoring unlabeled data and use labeled data only. This classic estimator $\widehat{\bm{\theta}}_{\textnormal{C}}$ solves for
$\dfrac{1}{n}\sum_{i=1}^{n}\bm{\psi}(y_{i},{\bf x}_{i};{\bm{\theta}})=0.$ (9)
It is always consistent, but can be of low efficiency when the labeled data is of a small size.
In this work, we aim to provide an estimator that is consistent and more efficient than the classic estimator, regardless of the quality of the $\widehat{f}$. In addition to the classic approach that takes average on labeled samples, we motivate our method with an augmented mean zero term that is indexed by a vector $\bm{\omega}$,
$\bm{\omega}\odot\left\{-\dfrac{1}{n}\sum_{i=1}^{n}\bm{\psi}(\widehat{f}({\bf z}_{i}),{\bf x}_{i};{\bm{\theta}})+\dfrac{1}{N}\sum_{i=n+1}^{n+N}\bm{\psi}(\widehat{f}({\bf z}_{i}),{\bf x}_{i};{\bm{\theta}})\right\},$ (10)
where $\odot$ is the element-wise multiplication. We refer to the vector $\bm{\omega}$ as a weighting vector as it controls the contribution of labeled outcomes and the arbitrary prediction $\widehat{f}$.
In light of this, we propose our POP-Inf estimator that solves the equation,
$\dfrac{1}{n}\sum_{i=1}^{n}\bm{\psi}(y_{i},{\bf x}_{i};{\bm{\theta}})+\widehat{\bm{\omega}}\odot\left\{-\dfrac{1}{n}\sum_{i=1}^{n}\bm{\psi}(\widehat{f}({\bf z}_{i}),{\bf x}_{i};{\bm{\theta}})+\dfrac{1}{N}\sum_{i=n+1}^{n+N}\bm{\psi}(\widehat{f}({\bf z}_{i}),{\bf x}_{i};{\bm{\theta}})\right\}={\bf 0},$ (11)
where $\widehat{\bm{\omega}}$ is an estimate of $\bm{\omega}$ aiming at minimizing the variance. We shall elaborate an optimal value in Section 3.1 and an estimation procedure in Section 3.2. For any weighting vector $\bm{\omega}$, the resulting estimator is denoted as $\widehat{\bm{\theta}}_{\textnormal{PopInf}}(\bm{\omega})$.
3.1 Asymptotic properties
We study the asymptotic properties of the proposed estimator $\widehat{\bm{\theta}}_{\textnormal{PopInf}}(\bm{\omega})$ with any weight $\bm{\omega}$. To establish its consistency and asymptotic normality, we introduce the following regularity conditions. Condition (C1) is a regularity assumption on the parameter of interest. Conditions (C2)-(C3) guarantee consistency and are weaker than assumptions typically used to guarantee uniform consistency (Newey and McFadden,, 1994). Condition (C4) is needed for asymptotic normality. All conditions are reasonable and standard (Van der Vaart,, 2000).
  • (C1)
    The parameter space ${\bm{\Theta}}$ is a compact subset of $\mathbb{R}^{q}$, containing the true parameter ${\bm{\theta}}_{0}$ as a unique solution to (1).
  • (C2)
    The function in (11) is continuous and the equation has a unique solution.
  • (C3)
[…]
  • (C4)
    There exists a function $\bm{\varphi}$ such that $\mathbb{E}[\sup_{{\bm{\theta}}\in\Theta}\bm{\varphi}^{2}]<\infty$ and $\mathbb{E}\lVert\bm{\psi}(Y,{\bf X};{\bm{\theta}}_{1})-\bm{\psi}(Y,{\bf X};{\bm{\theta}}_{2})\lVert_{2}^{2}\leq\bm{\varphi}\lVert{\bm{\theta}}_{1}-{\bm{\theta}}_{2}\rVert^{2}$ for any ${\bm{\theta}}_{1}$ and ${\bm{\theta}}_{2}$ in a neighborhood of ${\bm{\theta}}_{0}$. Furthermore, $\mathbb{E}\{\bm{\psi}(y,{\bf x};{\bm{\theta}})\}$ is differentiable at ${\bm{\theta}}_{0}$ with a nonsingular derivative matrix.
Theorem 1 (Asymptotic Normality).
Under Conditions (C1)-(C3), assuming $\frac{n}{N}\rightarrow\rho$ as $n\to\infty$ and $N\rightarrow\infty$, then the proposed estimator $\widehat{\bm{\theta}}_{\textnormal{PopInf}}(\bm{\omega})$ converges to ${\bm{\theta}}_{0}$ in probability. Assuming additionally Condition (C4), we have that as $n\to\infty$ and $N\to\infty$,
$\sqrt{n}(\widehat{\bm{\theta}}_{\textnormal{PopInf}}({\bm{\omega}})-{\bm{\theta}}_{0})\stackrel{{\scriptstyle D}}{{\rightarrow}}\mathcal{N}({\bf 0},{\bm{\Sigma}}(\bm{\omega})),$ (12)
where ${\bm{\Sigma}}(\bm{\omega})={\bf A}^{-1}{\bf V}(\bm{\omega}){\bf A}$, ${\bf A}=\mathbb{E}[\nabla_{{\bm{\theta}}}\bm{\psi}(Y,{\bf X};{\bm{\theta}}_{0})]$, ${\bf V}(\bm{\omega})={\bf M}_{1}+\bm{\omega}\bm{\omega}^{\rm T}\odot({\bf M}_{2}+\rho{\bf M}_{3})-2{\bf I}_{q}\bm{\omega}^{\rm T}\odot{\bf M}_{4}$, ${\bf M}_{1}=\textnormal{Var}_{l}[\bm{\psi}(Y,{\bf X};{\bm{\theta}}_{0})]$, ${\bf M}_{2}=\textnormal{Var}_{l}[\bm{\psi}(\widehat{f}({\bf Z}),{\bf X};{\bm{\theta}}_{0})]$, ${\bf M}_{3}=\textnormal{Var}_{u}[\bm{\psi}(\widehat{f}({\bf Z}),{\bf X};{\bm{\theta}}_{0})]$, ${\bf M}_{4}=\textnormal{Cov}_{l}[\bm{\psi}(Y,{\bf X};{\bm{\theta}}_{0}),\bm{\psi}(\widehat{f}({\bf Z}),{\bf X};{\bm{\theta}}_{0})]$.
Theorem 1 suggests that the proposed estimator $\widehat{\bm{\theta}}_{\textnormal{PopInf}}({\bm{\omega}})$ is asymptotically normal with any weight vector $\bm{\omega}$. With the pursuit of minimizing the asymptotic variance, we define the optimal weight $\bm{\omega}_{0}=(\omega_{01},\cdots,\omega_{0q})^{\rm T}$ such that each element $\omega_{0j}$ minimizes the variance of the corresponding element in the parameter,
$\omega_{0j}=\arg\min\limits_{\omega_{j}\in\mathbb{R}}{\bm{\Sigma}}_{j}{(\bm{\omega})}\text{ for }j=1,\cdots,q,$ (13)
where ${\bm{\Sigma}}_{j}(\bm{\omega})$ is the $j$-th diagonal element of the covariance matrix ${\bm{\Sigma}}{(\bm{\omega})}$. A data adaptive estimation procedure for $\bm{\omega}$ is provided in Section 3.2.
Remark 3 (Variance Comparison with the Classic Approach).
The asymptotic variance in Theorem 1 implies a direct comparison with the classic approach with labeled data only. Let $v_{j}(\omega_{j})$ be the $j$-th diagonal element of matrix ${\bf V}(\bm{\omega})$, then
$v_{j}(\omega_{j})=m_{1j}+(m_{2j}+\rho m_{3j})\omega_{j}^{2}-2m_{4j}\omega_{j},$ (14)
where $m_{1j}$, $m_{2j}$, $m_{3j}$ and $m_{4j}$ are the $j$-th diagonal element of matrices ${\bf M}_{1}$, ${\bf M}_{2}$, ${\bf M}_{3}$ and ${\bf M}_{4}$, respectively. The classic approach with $\omega_{j}=0$ corresponds to
$v_{j}(0)=m_{1j}.$ (15)
We compare it with the minimum of $v_{j}(\omega_{j})$ at two cases of $\rho$, the limit of $n/N$.
  • When $\rho\in[0,\infty)$, the minimum of $v_{j}(\omega_{j})$ is no greater than $v_{j}(0)$,
    $\min_{\omega_{j}}v_{j}(\omega_{j})=v_{j}(m_{4j}/(m_{2j}+\rho m_{3j}))$ $=m_{1j}-\dfrac{m_{4j}^{2}}{m_{2j}+\rho m_{3j}}\leq m_{1j},$ (16)
    suggesting the pursuit of a smaller variance. Note that $\rho\in[0,\infty)$ contains the special case $\rho=0$ which is common in the literature of semi-supervised learning.
  • When $\rho=\infty$, that is when $N=0$, we have $\min_{\omega_{j}}v_{j}(\omega_{j})=v_{j}(0)$. This suits our intuition that when all data are labeled, the optimal choice is asymptotically as efficient as the classic approach.
Remark 4 (Data-adaptive Variance Reduction).
In (14), we notice that the last term is $-2m_{4j}\omega_{j}$. Recall that $m_{4j}$ is the $j$-th diagonal entry of the covariance matrix between true outcomes and predicted outcomes on the labeled data, it reflects the prediction accuracy: higher $m_{4j}$ implies higher prediction accuracy. If $\omega_{j}$ is fixed and nonzero, then $-2m_{4j}\omega_{j}$ becomes smaller as $m_{4j}$ increases and vice versa, indicating that a better prediction algorithm yields a smaller variance. While this observation should hold in general, our proposed POP-Inf method allows a data-adaptive choice of $\omega_{j}$ so that it amplifies variance reduction when the prediction performs well, and avoids variance inflation otherwise.
Remark 5 (Element-wise Variance Reduction).
The optimal choice weight (19) is defined for every element of the parameter so that each component is targeted to have a reduced variance. Under cases where only a subset of the parameter is of inferential interest, it’s possible to only assign nontrivial weights to corresponding elements. For example in a regression model, if one only aims to make inference on the coefficient of a particular covariate, then the other $\omega_{j}$’s can be set to zero.
The asymptotic normality established in Theorem 1 enables variance estimation and statistical inference. Corollary 1 illustrates the construction of a confidence interval, where a consistent estimate $\widehat{\bm{\Sigma}}{(\bm{\omega})}$ of the covariance matrix can be easily obtained by the sample counterpart of ${\bm{\Sigma}}{(\bm{\omega})}$.
Corollary 1 (Confidence Interval).
Suppose $\widehat{\bm{\Sigma}}{(\bm{\omega})}\stackrel{{\scriptstyle P}}{{\rightarrow}}{\bm{\Sigma}}{(\bm{\omega})}$. Given confidence level $1-\alpha\in(0,1)$, construct an interval for the $j$-th element of the parameter, $\theta_{0,j}$, as
${\cal C}_{\alpha,j}=\left\{\theta:\theta\in\left(\widehat{\theta}_{\textnormal{PopInf}}({\bm{\omega}})_{j}-z_{\alpha/2}\{\widehat{\bm{\Sigma}}_{j}{(\bm{\omega})}/n\}^{-1/2},\widehat{\theta}_{\textnormal{PopInf}}({\bm{\omega}})_{j}+z_{\alpha/2}\{\widehat{\bm{\Sigma}}_{j}{(\bm{\omega})}/n\}^{-1/2}\right)\right\},$ (17)
where $\widehat{\theta}_{j}$ is the $j$-th element of $\widehat{\bm{\theta}}_{\textnormal{PopInf}}$ and $z_{\alpha/2}$ is the upper $\alpha/2$ quantile of the standard normal distribution. Then
$\lim_{N\in\infty}\mathbb{P}(\theta_{0,j}\in{\cal C}_{\alpha,j})\geq 1-\alpha,\ j=1\cdots,q.$ (18)
Hence, ${\cal C}_{\alpha,j}$ is a confidence interval for $\theta_{0,j}$ at level $1-\alpha$.
3.2 Estimation procedure
In practice, we propose to estimate the weight with $\widehat{\bm{\omega}}$ that minimizes each element of the sample counterpart of the covariance matrix. That is,
$\widehat{\omega}_{j}=\arg\min\limits_{\omega_{j}\in{\bf R}}\widehat{\bm{\Sigma}}_{j}{(\bm{\omega})}\text{ for }j=1,\cdots,q,$ (19)
where $\widehat{\bm{\Sigma}}_{j}{(\bm{\omega})}$ is the $j$-th diagonal element of the sample analog of ${\bm{\Sigma}}{(\bm{\omega})}$.
The following Algorithm 1 details an adaptive estimation procedure that updates the estimates for both the weight and the parameter of interest at each iteration. In the end, after obtaining the estimator and the corresponding weight, one can construct a confidence interval as stated in Corollary 1. Note that this algorithm is applicable to any parameter defined in (1). It can be simplified under certain scenarios. As illustrated in Section 2.1, no iterations are needed to estimate the outcome. In that example, one can consistently estimate the weight directly from data and build an estimation equation for estimating the outcome mean.
Algorithm 1 POP-Inf Estimation
1:Data ${\cal L}\cup{\cal U}$, predictor $\widehat{f}$, $\bm{\omega}^{0}={\bf 0}$, error rate $\alpha\in(0,1)$, a step size $\eta$ for gradient descent, tolerance $\delta$, and the maximum of iterations $K$.
2:Obtain an initial estimate ${\bm{\theta}}^{0}$ by solving $\dfrac{1}{n}\sum_{i=1}^{n}\bm{\psi}(y_{i},{\bf x}_{i};{\bm{\theta}})={\bf 0}$.
3:for $k$ in $1,\cdots\cdots,K$ do
4:     Update $\widehat{\bm{\omega}}^{k}$ with each element $\widehat{\omega}_{j}^{k}=\mbox{argmin}_{\omega_{j}\in\mathbb{R}}\widehat{\bm{\Sigma}}_{j}(\widehat{\bm{\omega}}^{k-1})\text{ for }j=1,\cdots,q,$.
5:     Update $\widehat{\bm{\theta}}^{k}$ by solving (11) with $\bm{\omega}^{k-1}$.
6:until convergence criterion $\|\widehat{\bm{\theta}}^{k}-\widehat{\bm{\theta}}^{k-1}\|<\delta$ satisfied.
7:POP-Inf weight $\widehat{\bm{\omega}}$ and the estimator $\widehat{\bm{\theta}}_{\textnormal{PopInf}}(\widehat{\bm{\omega}})$.
3.3 Examples: linear and logistic regression
Here, we give two examples to provide intuitions for POP-Inf estimation procedure. We start with estimating coefficients in a linear regression model where the resulting estimator has a closed form solution.
Example 2 (Linear Regression).
We consider the ordinary least squares problem, where $\bm{\psi}(y,{\bf x};{\bm{\theta}})={\bf x}(y-{\bf x}^{\rm T}{\bm{\theta}})$. Therefore, the estimand is ${\bm{\theta}}_{0}=\mathbb{E}[{\bf X}{\bf X}^{\rm T}]^{-1}\mathbb{E}[{\bf X}Y]$. From equation (11), the POP-Inf estimator should be the solution to the equation
$\dfrac{1}{n}\sum_{i=1}^{n}{\bf x}_{i}(y_{i}-{\bf x}_{i}^{\rm T}{\bm{\theta}})+\dfrac{1}{N}\sum_{i=n+1}^{n+N}\bm{\omega}\odot{{\bf x}}_{i}\{\widehat{f}({\bf z}_{i})-{\bf x}_{i}^{\rm T}{\bm{\theta}}\}-\dfrac{1}{n}\sum_{i=1}^{n}\bm{\omega}\odot{\bf x}_{i}\{\widehat{f}({\bf z}_{i})-{\bf x}_{i}^{\rm T}{\bm{\theta}}\}={\bf 0},$ (20)
yielding a closed-form solution for POP-Inf linear regression
$\widehat{\theta}_{\textnormal{PopInf}}({\widehat{\bm{\omega}}})$ =$=\left\{\dfrac{1}{n}\sum_{i=1}^{n}{\bf x}_{i}{\bf x}_{i}^{\rm T}+\widehat{\bm{\omega}}\odot\left[\dfrac{1}{N}\sum_{i=n+1}^{n+N}{{\bf x}}_{i}{\bf x}_{i}^{\rm T}-\dfrac{1}{n}\sum_{i=1}^{n}{\bf x}_{i}{\bf x}_{i}^{\rm T}\right]\right\}^{-1}\cdot\left\{\dfrac{1}{n}\sum_{i=1}^{n}{\bf x}_{i}y_{i}+\widehat{\bm{\omega}}\odot\left[\dfrac{1}{N}\sum_{i=n+1}^{n+N}{\bf x}_{i}{\widehat{f}}({\bf z}_{i})-\dfrac{1}{n}\sum_{i=1}^{n}{\bf x}_{i}\widehat{f}({\bf z}_{i})\right]\right\},$ (21)
where $\widehat{\bm{\omega}}$ is a consistent estimate of optimal $\bm{\omega}_{0}$, which can be obtained by Algorithm 1.
The second example is logistic regression, which does not have a closed-form solution.
Example 3 (Logistic Regression).
For logistic regression, $\bm{\psi}(y,{\bf x};{\bm{\theta}})=-{\bf x}y+{\bf x}\gamma_{{\bm{\theta}}}({\bf x})$, where $\gamma_{{\bm{\theta}}}({\bf x})=1/\left\{1+\exp\left(-{\bf x}^{\rm T}{\bm{\theta}}\right)\right\}$. Using estimation equation (11), the POP-Inf estimator $\widehat{\bm{\theta}}_{\textnormal{PopInf}}({\widehat{\bm{\omega}}})$ isthe solution to the equation
$\dfrac{1}{n}\sum_{i=1}^{n}\left\{-{\bf x}_{i}y_{i}+{\bf x}_{i}\gamma_{{\bm{\theta}}}({\bf x}_{i})\right\}+\bm{\omega}\odot\left\{\dfrac{1}{N}\sum_{i=n+1}^{n+N}[-{{\bf x}}_{i}\widehat{f}({\bf z}_{i})+{{\bf x}}_{i}\gamma_{{\bm{\theta}}}({\bf x}_{i})]-\dfrac{1}{n}\sum_{i=1}^{n}[-{\bf x}_{i}\widehat{f}({\bf z}_{i})+{\bf x}_{i}\gamma_{{\bm{\theta}}}({\bf x}_{i})]\right\}={\bf 0}.$ (22)
This does not yield a closed form solution. To obtain an estimator, one can adopt Algorithm 1.
4 Relationship with Efficient Influence Function
To connect our method with the literature of semiparametric efficiency (Bickel et al.,, 1993; Tsiatis,, 2006), in this section, we state the efficient influence function (EIF) and the semiparametric efficiency bound for the estimation problem (1) (Proposition 1) and connect it with our proposed POP-Inf estimator.
To ease notation, we introduce a random variable $R$ to indicate the labeling; $r_{1}=\cdots=r_{n}=1$ and $r_{n+1}=\cdots=r_{n+N}=1$. In this section, suppose $\pi\equiv\hbox{pr}(R=1)=n/(n+N)$ is a fixed constant between zero and one. Note that the introduction of $R$ is for notational simplicity and we make no inference on it.
Proposition 1 (Efficient Influence Function).
The EIF for estimating ${\bm{\theta}}$ is
$\bm{\phi}(y_{i},{\bf x}_{i},{\bm{\theta}})=\dfrac{r}{\pi}{\bf A}^{-1}[\bm{\psi}(y,{\bf x};{\bm{\theta}})-\mathbb{E}\{\bm{\psi}(Y,{\bf x};{\bm{\theta}})\mid{\bf x},{\bf z}\}]+{\bf A}^{-1}\mathbb{E}\{\bm{\psi}(Y,{\bf x};{\bm{\theta}})\mid{\bf x},{\bf z}\}.$ (23)
Guided by Theorem 1, an EIF-based estimator would be the solution to the empirical equation,
${\bf 0}=\dfrac{1}{n+N}\sum_{i=1}^{n+N}\bm{\phi}(y_{i},{\bf x}_{i},\widehat{\bm{\theta}}_{\textnormal{EIF}}).$ (24)
As stated in Corollary 2, this EIF-basd estimator is asymptotically normal and achieves the semiparametric efficiency bound.
Corollary 2 (Semiparametric Efficiency Bound).
Under Conditions (C1) and (C3), assume $\mathbb{E}\{\sup_{{\bm{\theta}}\in{\bm{\Theta}}}\bm{\phi}(Y,{\bf X},{\bm{\theta}})\}<\infty$ and equation (24) has a unique solution, then as $n\to\infty$ and $N\to\infty$,
$\sqrt{n}(\widehat{\bm{\theta}}_{\textnormal{EIF}}-{\bm{\theta}}_{0})\stackrel{{\scriptstyle D}}{{\rightarrow}}\mathcal{N}({\bf 0},\mathbb{E}\{\bm{\phi}(Y,{\bf X},{\bm{\theta}}_{0})\bm{\phi}(Y,{\bf X},{\bm{\theta}}_{0})^{\rm T}\pi\}).$ (25)
If the nuisance function that appears in the EIF, in particular $\mathbb{E}(Y\mid{\bf x},{\bf z})$ here, can be correctly specified or well approximated, then an EIF-based estimator is feasible in practice, and it will achieve the semiparametric efficiency bound. However, this is nearly impossible in reality with limited knowledge of the data generation process or a small computational budget. Instead, it’s typical to construct an estimator with nuisance functions plugged in as their estimates. We refer to such estimators as “$*{}^{*}$-based”, because the influence function of the resulting estimator, denoted by $\widehat{\bm{\theta}}_{\textnormal{EIF}^{*}}$, may or may not be the efficient influence function.
In our problem, the nuisance function $\mathbb{E}(Y\mid{\bf x},{\bf z})$ can be estimated with $\widehat{f}({\bf z})$, which can be far from satisfactory even if the covariate of interest ${\bf x}$ is a subset of the predictor ${\bf z}$. Upon simple calculation, we can write the resulting $\widehat{\bm{\theta}}_{\textnormal{EIF}^{*}}$ as the solution to
$\dfrac{1}{n}\sum_{i=1}^{n}\bm{A}^{-1}\bm{\psi}(y_{i},{\bf x}_{i};{\bm{\theta}})+(1-\pi){\bf A}^{-1}\left\{-\dfrac{1}{n}\sum_{i=1}^{n}\bm{\psi}(\widehat{f}({\bf z}_{i}),{\bf x}_{i};{\bm{\theta}})+\dfrac{1}{N}\sum_{i=n+1}^{n+N}\bm{\psi}(\widehat{f}({\bf z}_{i}),{\bf x}_{i};{\bm{\theta}})\right\}={\bf 0}.$ (26)
Similar to the equation (11) that the POP-Inf estimator solves, the $*{}^{*}$-based estimator (26) solves a equation with two parts: one term that can produce a consistent estimator with labeled data only, and the other term utilizes predictions $\widehat{f}$ that is always mean zero. If one multiples both sides of (11) with matrix ${\bf A}^{-1}$, then (26) corresponds to that equation weight equals to a vector of $1-\pi$. In estimating the outcome mean (Section 2.1) where ${\bf A}$ is essentially one, the $*{}^{*}$-based estimator is the POP-Inf estimator with weight being $1-\pi=N/(n+N)$. The difference between (11) and (26) clarifies one of the main distinctions between the proposed POP-Inf estimator and an EIF guided estimator. The former seeks for variance reduction in a data-adaptive way, while the latter attains minimum variance at correct specifications. In real-world situations, the proposed POP-Inf estimator can be much more efficient in that it utilizes information in a data-adaptive fashion. Numeric comparisons are provided in Section 5.
5 Simulations
To demonstrate our method and corroborate our theoretical findings, we estimate outcome mean and linear regression coefficients on simulated datasets with the POP-Inf estimator and its competitors. In a variety of simulated datasets, all methods lead to satisfactory coverage. Moreover, the proposed POP-Inf estimator yields the smallest variances. We detail the data generation and implementation in Section 5.1 and present results in Section 5.2.
5.1 Settings
We generate a dataset with $n\in\{250,500\}$ labeled samples and $N=20n$ unlabeled samples, identically and independently drawn from the following data generating process. For unit $i$, we generate ${\bf X}_{i}=[X_{1i},X_{2i}]^{\rm T}$ from a multivariate normal distribution with mean zero and an identity covariance matrix. Without loss of generality, we set $r_{1}=\cdots=r_{n}=1$ and generate their outcome
$Y_{i}=5\gamma\left(X_{1i}+X_{1i}+X_{1i}^{2}+X_{2i}^{2}+X_{1i}X_{2i}\right)/\sqrt{6}+\varepsilon_{i},$ (27)
where $\varepsilon_{i}\sim\mathcal{N}(0,5\sqrt{1-\gamma^{2}})$ independently and $\gamma$ varies between zero and one by a grid of $0.02$. The parameter $\gamma$ controls the data generation and it quantifies the signal to noise ratio. Roughly speaking, the higher $\gamma$ is, the easier it would be to predict $Y$, and henceforth the higher efficiency one would expect for any estimator that utilizes unlabeled data.
For the simplicity of this simulation, we consider ${\bf Z}={\bf X}$. To build the ML prediction algorithm $\widehat{f}$, we generate an independent training set of size $500$ with the same data generating process, on which we train a random forest with R package randomForest (Breiman,, 2001). Arguments have been set to default .
We consider two estimands, the outcome mean and the linear regression coefficient of regressing $Y$ onto $X_{1}$ with independent and identically distributed normal errors. For each estimand, we construct four estimators: the classic estimator, the PP estimator (Angelopoulos et al., 2023a, ), the $*{}^{*}$-based estimator, and the proposed POP-Inf estimator. Simulation results are based on $1000$ repetitions. At each repetition, we calculate standard errors for all estimators based on Theorem 1 and Theorem 1.
5.2 Results
As shown in the left panels in Figure 1, all methods lead to a coverage rate at around the nominal level $95\%$ when $n=250$, suggesting that they are all unbiased with valid standard error estimates. To compare the efficiency of an estimator compared with the classic approach,, we calculate the ratios of standard error estimates between one estimator and the classic approach. When the ratio is smaller than one, it suggests that the corresponding estimator is more efficient than the classic approach and vice versa. The right panels of Figure 1 plot these ratios against $\gamma\in[0,0.5]$ for POP-Inf, PP, and $*{}^{*}$-based estimators for both estimands at $n=250$. These estimators behave similarly for the two estimands.
When $\gamma$ is close to zero, the POP-Inf estimator is slightly more efficient than the classic approach, while the other two estimators are much less efficient. As $\gamma$ increases to $0.5$, reflecting that $\widehat{f}$ improves, all ratios drop. The POP-Inf estimator always has the lowest ratio, meaning its efficiency improvement against other estimators. When $\gamma\in[0.5,1]$, as can be seen in Figure 2, the POP-Inf estimator still has the smallest standard error, whereas the other two estimators have comparable performance. This suggests that when the prediction performs well, the proposed POP-Inf estimator is still a better choice. Similar patterns can be found for $n=500$ (Figure 3). In addition, Figure 4 shows boxplots of biases of the POP-Inf estimator across $1000$ repetitions. As the sample size increases, the boxes become narrower and hence the variance becomes smaller. This corroborates the root $n$ consistency.
Figure 2: Relative efficiency in simulations with $\gamma\in[0.5,1]$ and $n=250$. In each panel, the x-axis coordinates the data generating parameter $\gamma$ detailed in Section 5. Left panels plot the ratios between averaged standard errors of a competing method and the classic approach, with dashed horizontal line intersects $1$ at the y-axis. A ratio lower than one represents a smaller standard error than the classic approach.
Figure 3: Relative efficiency and coverage rates in simulations with $\gamma\in[0,1]$ and $n=500$. In each panel, the x-axis coordinates the data generating parameter $\gamma$ detailed in Section 5. Left panels plot the ratios between averaged standard errors of a competing method and the classic approach, with dashed horizontal line intersects $1$ at the y-axis. A ratio lower than one represents a smaller standard error than the classic approach.
Figure 4: Bias of the POP-Inf estimator in simulations for (red) $n=250$ and (purple) $n=500$. Panels are faceted by the estimand and coordinated by $\gamma\in[0,1]$ on the x-axis.
6 Data Applications
We used the POP-Inf method to assess the effect of biological sex on gene expression across 44 human tissues using GTEx data described in Example 1. We aim to use a ML algorithm called hypergraph factorization (HYFA) to impute the missing gene expression (Viñas et al.,, 2023) in the uncollected tissue. The imputed gene expression is then used in POP-Inf to increase the statistical power to identify sex-biased genes. We compared POP-Inf with classic, PP, and $*{}^{*}$-based approaches.
In particular, we processed the GTEx-v8 data following the GTEx pipeline (GTEx Consortium et al.,, 2015). We then used HYFA to predict transcript levels for each uncollected tissue for each individual in the data, resulting in a gene expression dataset of $834$ individuals across $44$ tissues. HYFA is a parameter-efficient graph representation learning approach for multi-tissue prediction of gene expression. We used cross-validation to predict transcript levels in the labeled data to avoid over-fitting. Prediction accuracy is measured by the correlation between measured and imputed gene expression in the labeled data. We then used linear regression to assess the effect of sex on gene expression while controlling for technical factors such as surrogate variables, age, RNA integrity number, and total ischemic time for a sample.
Results are shown in Figure 5. Each dot represents the result of the inference for one gene in one tissue in Figure 5 (a)-(f). The point estimates of all method are close to the classic approach (Figure 5 (a)-(c)). However, both PP and $*{}^{*}$-based are less efficient than the classic approach when the prediction accuracy is low. In contrast, the POP-Inf estimator is always no less efficient than the classic approach, regardless of the prediction accuracy (Figure 5 (d)-(f)). Moreover, POP-Inf identifies more sex-biased genes than other approaches, demonstrating its improved efficiency over alternatives (Figure 5 (g)).
In conclusion, the real data results are consistent with our theoretical and simulation results, demonstrating that our method POP-Inf is more attractive in real-world post-prediction inference.
Figure 5: Comparison of POP-Inf, classic, PP, and $*{}^{*}$-based approaches in identifying sex-biased genes using GTEx data. Each panel illustrates a different aspect of comparison on the y- and x- axes: point estimates between the (a) classic and PP approaches; (b) classic and $*{}^{*}$-based approaches; (c) classic and POP-Inf approaches; estimated standard errors of point estimates between the (d) classic and PP approaches; (e) classic and $*{}^{*}$-based approaches; (f) classic and POP-Inf approaches; (g) number of sex-biased genes identified by each of the four approaches. The dashed line represent $y=x$ in (a)-(c) and $y=1$ in (d)-(f).
7 Discussion
We have provided a simple yet powerful method, the POP-Inf, to improve the efficiency of using unlabeled data with arbitrary ML predictions. To enable inference, we establish the consistency and asymptotic normality of the proposed estimator and prove its superiority over baseline methods. In extensive simulations and real data application, we demonstrate the superiority of the POP-Inf estimator over alternatives.
After finishing this manuscript, we became aware of an independent work by Angelopoulos et al., 2023b that was preprinted in the same month as ours. This work is similar to POP-Inf in using weights to guarantee an efficiency improvement with arbitrary ML predictions. Our development of POP-Inf was conducted independently and without prior knowledge of their work. Moreover, our approach differs in several key aspects. First, we incorporate a weight vector in the the estimating equation, while Angelopoulos et al., 2023b add a scalar weight to the loss function. Therefore, our approach optimizes the weight in a larger set of estimators. Second, we targets an element-wise variance reduction of the (vector) parameter of interest, while Angelopoulos et al., 2023b improve efficiency by reducing the trace of the covariance matrix. Therefore, our method has guaranteed efficiency gain over the classic method for each parameter. This is important in practice, because many statistical analyses have a specific subvector of the parameter garnering scientific attention, whereas other parameters are considered as nuisances (GTEx Consortium,, 2020; Uffelmann et al.,, 2021). Minimizing the trace of the covariance matrix for all elements including those for the nuisance parameters does not guarantee element-wise variance reduction, which can lead to inflated variances for the parameters of interest compared to classic methods. Third, we connect our estimator to the efficient influence function, which is another powerful tool in the pursuit of efficiency. We have demonstrated in numeric comparisons that the POP-Inf estimator is more practical than an $*{}^{*}$-based estimator when a good estimate of the nuisance function cannot be guaranteed.
This work fits into a variety of scientific problems and unlocks several directions of future research. One can apply the same principal to more complicated data structure and missing mechanisms, for example in cases where a multi-dimensional outcome or covariate is subject to missing. It would be an interesting future work to carefully discuss relationships between ${\bf X}$ and ${\bf Z}$ and and how they may affect the magnitude of the efficiency improvement.
Acknowledgments
We gratefully acknowledge research support from the University of Wisconsin-Madison Office of the Chancellor and the Vice Chancellor for Research and Graduate Education with funding from the Wisconsin Alumni Research Foundation (WARF). We thank the participants of the UW-Madison Statistics Graduate Student Association Student Seminar and the Lu Group Lab Meeting for many useful comments and questions.

### 04 Feb 2024 v2

@@
 Further, we establish the asymptotic properties of the POP-Inf estimator and describe the estimation procedure.
-We build the connection with the semiparametric efficiency theory in Section 4.
+We build the connection with the semiparametric efficiency theory in Section 4 and describe extensions in Section 5.
 Through extensive simulations (Section 6) and real data analysis (Section 7), we validate our theoretical claims and demonstrate the practical utility and effectiveness of the POP-Inf estimator.
@@
 The resulting protocol can benefit researches where training a prediction algorithm is unrealistic (e.g., training ChatGPT), and provide guidelines for scientists who are not equipped with computational resources or knowledge about building prediction algorithms.
-Remark 1.
-(The consideration of ${\bf X}$ and ${\bf Z}$)
+Remark 1 (The consideration of ${\bf X}$ and ${\bf Z}$).
 We deliberately split the covariate into two (possibly overlapping) parts, ${\bf Z}$ and ${\bf X}$, where the former is predictive of the outcome $Y$ and the latter appears in the estimating equation (1).
@@
 We provide the intuition of our approach using mean estimation.
-Consider $U(y,{\bf x};{\bm{\theta}})=y-\theta$.
+Consider $\psi(y,{\bf x};{\bm{\theta}})=y-\theta$.
 The resulting estimand is $\theta=\mathbb{E}(Y)$, the outcome mean.
@@
 To increase efficiency with the ML-predicted outcomes in unlabeled data, we propose to introduce another unbiased estimator of zero with a weight scalar $\omega$,
-$\widehat{\theta}_{\textnormal{PopInf}}(\omega)=\dfrac{1}{n}\sum_{i=1}^{n}y_{i}+\omega\left[\dfrac{1}{N}\sum_{i=n+1}^{n+N}\widehat{f}({\bf z}_{i})-\dfrac{1}{n}\sum_{i=1}^{n}\widehat{f}({\bf z}_{i})\right]$ (3)
+$\widehat{\theta}_{\textnormal{POP-Inf}}(\omega)=\dfrac{1}{n}\sum_{i=1}^{n}y_{i}+\omega\left[\dfrac{1}{N}\sum_{i=n+1}^{n+N}\widehat{f}({\bf z}_{i})-\dfrac{1}{n}\sum_{i=1}^{n}\widehat{f}({\bf z}_{i})\right]$ (3)
 We refer to this estimator as the POP-Inf estimator.
@@
 In addition, Wald-type confidence intervals can be constructed using the following asymptotic variance,
-$\underbrace{\frac{\textnormal{Var}_{l}(Y)}{n}+\underbrace{\left(\omega^{2}\frac{\textnormal{Var}_{l}[\widehat{f}]}{n}-2\omega\frac{\textnormal{Cov}_{l}[Y,\widehat{f}]}{n}+\omega^{2}\frac{\textnormal{Var}_{u}[\widehat{f}]}{N}\right)}_{\text{Additional terms}}}_{\text{{POP-Inf}\leavevmode\nobreak\ interval}},$ (6)
+$\underbrace{\frac{\textnormal{Var}_{l}(Y)}{n}+\underbrace{\left(\omega^{2}\frac{\textnormal{Var}_{l}[\widehat{f}]}{n}-2\omega\frac{\textnormal{Cov}_{l}[Y,\widehat{f}]}{n}+\omega^{2}\frac{\textnormal{Var}_{u}[\widehat{f}]}{N}\right)}_{\text{Additional terms}}}_{\text{Variance of {POP-Inf}\leavevmode\nobreak\ estimator}},$ (6)
 where $\textnormal{Var}_{l}[\cdot]$ and $\textnormal{Cov}_{l}[\cdot,\cdot]$ represent the variance and covariance on the labeled, respectively, and $\textnormal{Var}_{u}[\cdot]$ is the variance on the unlabeled.
@@
 $q(\omega)=\omega^{2}\left\{\frac{\textnormal{Var}_{l}[\widehat{f}]}{n}+\frac{\textnormal{Var}_{u}[\widehat{f}]}{N}\right\}-2\omega\frac{\textnormal{Cov}_{l}[Y,\widehat{f}]}{n},$ (7)
-which achieves its minimum at $\omega_{0}=\textnormal{Cov}_{l}[Y,\widehat{f}({\bf z})]/\{\textnormal{Var}_{l}[\widehat{f}]+n\textnormal{Var}_{l}[\widehat{f}]/N\}$ with minimum value
+which achieves its minimum at $\omega_{0}=\textnormal{Cov}_{l}[Y,\widehat{f}({\bf z})]/\{\textnormal{Var}_{l}[\widehat{f}]+n\textnormal{Var}_{u}[\widehat{f}]/N\}$ with minimum value
 $q(\omega_{0})=\dfrac{-\textnormal{Cov}_{l}[Y,\widehat{f}]}{n\textnormal{Var}_{l}[\widehat{f}]+n^{2}\textnormal{Var}_{u}[\widehat{f}]/N}\leq 0,$ (8)
@@
 In light of this, we propose our POP-Inf estimator that solves the equation,
-$\dfrac{1}{n}\sum_{i=1}^{n}\bm{\psi}(y_{i},{\bf x}_{i};{\bm{\theta}})+\widehat{\bm{\omega}}\odot\left\{-\dfrac{1}{n}\sum_{i=1}^{n}\bm{\psi}(\widehat{f}({\bf z}_{i}),{\bf x}_{i};{\bm{\theta}})+\dfrac{1}{N}\sum_{i=n+1}^{n+N}\bm{\psi}(\widehat{f}({\bf z}_{i}),{\bf x}_{i};{\bm{\theta}})\right\}={\bf 0},$ (11)
-where $\widehat{\bm{\omega}}$ is an estimate of $\bm{\omega}$ aiming at minimizing the variance.
+$\dfrac{1}{n}\sum_{i=1}^{n}\bm{\psi}(y_{i},{\bf x}_{i};{\bm{\theta}})+\bm{\omega}\odot\left\{-\dfrac{1}{n}\sum_{i=1}^{n}\bm{\psi}(\widehat{f}({\bf z}_{i}),{\bf x}_{i};{\bm{\theta}})+\dfrac{1}{N}\sum_{i=n+1}^{n+N}\bm{\psi}(\widehat{f}({\bf z}_{i}),{\bf x}_{i};{\bm{\theta}})\right\}={\bf 0},$ (11)
+where $\bm{\omega}$ can be replaced with an estimate $\widehat{\bm{\omega}}$ aiming at minimizing the variance.
 We shall elaborate an optimal value in Section 3.1 and an estimation procedure in Section 3.2.
@@
 Condition (C1) is a regularity assumption on the parameter of interest.
-Conditions (C2)-(C3) guarantee consistency and are weaker than assumptions typically used to guarantee uniform consistency (Newey and McFadden,, 1994).
+Conditions (C2)-(C3) guarantee consistency.
 Condition (C4) is needed for asymptotic normality.
@@
 Assuming additionally Condition (C4), we have that as $n\to\infty$ and $N\to\infty$,
-$\sqrt{n}(\widehat{\bm{\theta}}_{\textnormal{PopInf}}({\bm{\omega}})-{\bm{\theta}}_{0})\stackrel{{\scriptstyle D}}{{\rightarrow}}\mathcal{N}({\bf 0},{\bm{\Sigma}}(\bm{\omega})),$ (12)
-where ${\bm{\Sigma}}(\bm{\omega})={\bf A}^{-1}{\bf V}(\bm{\omega}){\bf A}$, ${\bf A}=\mathbb{E}[\nabla_{{\bm{\theta}}}\bm{\psi}(Y,{\bf X};{\bm{\theta}}_{0})]$, ${\bf V}(\bm{\omega})={\bf M}_{1}+\bm{\omega}\bm{\omega}^{\rm T}\odot({\bf M}_{2}+\rho{\bf M}_{3})-2{\bf I}_{q}\bm{\omega}^{\rm T}\odot{\bf M}_{4}$, ${\bf M}_{1}=\textnormal{Var}_{l}[\bm{\psi}(Y,{\bf X};{\bm{\theta}}_{0})]$, ${\bf M}_{2}=\textnormal{Var}_{l}[\bm{\psi}(\widehat{f}({\bf Z}),{\bf X};{\bm{\theta}}_{0})]$, ${\bf M}_{3}=\textnormal{Var}_{u}[\bm{\psi}(\widehat{f}({\bf Z}),{\bf X};{\bm{\theta}}_{0})]$, ${\bf M}_{4}=\textnormal{Cov}_{l}[\bm{\psi}(Y,{\bf X};{\bm{\theta}}_{0}),\bm{\psi}(\widehat{f}({\bf Z}),{\bf X};{\bm{\theta}}_{0})]$.
+$\sqrt{n}(\widehat{\bm{\theta}}_{\textnormal{POP-Inf}}({\bm{\omega}})-{\bm{\theta}}_{0})\stackrel{{\scriptstyle D}}{{\rightarrow}}\mathcal{N}({\bf 0},{\bm{\Sigma}}(\bm{\omega})),$ (12)
+where ${\bm{\Sigma}}(\bm{\omega})={\bf A}^{-1}{\bf V}(\bm{\omega}){\bf A}^{-1}$, ${\bf A}=\mathbb{E}[\partial\bm{\psi}(Y,{\bf X};{\bm{\theta}}_{0})/\partial{\bm{\theta}}]$, ${\bf V}(\bm{\omega})={\bf M}_{1}+\bm{\omega}\bm{\omega}^{\rm T}\odot({\bf M}_{2}+\rho{\bf M}_{3})-2{\bf I}_{q}\bm{\omega}^{\rm T}\odot{\bf M}_{4}$, ${\bf M}_{1}=\textnormal{Var}_{l}[\bm{\psi}(Y,{\bf X};{\bm{\theta}}_{0})]$, ${\bf M}_{2}=\textnormal{Var}_{l}[\bm{\psi}(\widehat{f}({\bf Z}),{\bf X};{\bm{\theta}}_{0})]$, ${\bf M}_{3}=\textnormal{Var}_{u}[\bm{\psi}(\widehat{f}({\bf Z}),{\bf X};{\bm{\theta}}_{0})]$, ${\bf M}_{4}=\textnormal{Cov}_{l}[\bm{\psi}(Y,{\bf X};{\bm{\theta}}_{0}),\bm{\psi}(\widehat{f}({\bf Z}),{\bf X};{\bm{\theta}}_{0})]$.
+Proof is contained in Appendix B.1.
 Theorem 1 suggests that the proposed estimator $\widehat{\bm{\theta}}_{\textnormal{POPInf}}({\bm{\omega}})$ is asymptotically normal with any weight vector $\bm{\omega}$.
@@
 Given confidence level $1-\alpha\in(0,1)$, construct an interval for the $j$-th element of the parameter, $\theta_{0,j}$, as
-${\cal C}_{\alpha,j}=\left\{\theta:\theta\in\left(\widehat{\theta}_{\textnormal{PopInf}}({\bm{\omega}})_{j}-z_{\alpha/2}\{\widehat{\bm{\Sigma}}_{j}{(\bm{\omega})}/n\}^{-1/2},\widehat{\theta}_{\textnormal{PopInf}}({\bm{\omega}})_{j}+z_{\alpha/2}\{\widehat{\bm{\Sigma}}_{j}{(\bm{\omega})}/n\}^{-1/2}\right)\right\},$ (17)
+${\cal C}_{\alpha,j}=\left\{\theta:\theta\in\left(\widehat{\theta}_{\textnormal{POP-Inf}}({\bm{\omega}})_{j}-z_{\alpha/2}\{\widehat{\bm{\Sigma}}_{j}{(\bm{\omega})}/n\}^{-1/2},\widehat{\theta}_{\textnormal{POP-Inf}}({\bm{\omega}})_{j}+z_{\alpha/2}\{\widehat{\bm{\Sigma}}_{j}{(\bm{\omega})}/n\}^{-1/2}\right)\right\},$ (17)
 where $\widehat{\theta}_{j}$ is the $j$-th element of $\widehat{\bm{\theta}}_{\textnormal{POPInf}}$ and $z_{\alpha/2}$ is the upper $\alpha/2$ quantile of the standard normal distribution. Then
@@
 3:for $k$ in $1,\cdots\cdots,K$ do
-4:     Update $\widehat{\bm{\omega}}^{k}$ with each element $\widehat{\omega}_{j}^{k}=\mbox{argmin}_{\omega_{j}\in\mathbb{R}}\widehat{\bm{\Sigma}}_{j}(\widehat{\bm{\omega}}^{k-1})\text{ for }j=1,\cdots,q,$.
+4:     Update $\widehat{\bm{\omega}}^{k}$ with each element $\widehat{\omega}_{j}^{k}=\mbox{argmin}_{\omega_{j}\in\mathbb{R}}\widehat{\bm{\Sigma}}_{j}(\widehat{\bm{\omega}}^{k-1})\text{ for }j=1,\cdots,q$.
 5:     Update $\widehat{\bm{\theta}}^{k}$ by solving (11) with $\bm{\omega}^{k-1}$.
 6:until convergence criterion $\|\widehat{\bm{\theta}}^{k}-\widehat{\bm{\theta}}^{k-1}\|<\delta$ satisfied.
-7:POP-Inf weight $\widehat{\bm{\omega}}$ and the estimator $\widehat{\bm{\theta}}_{\textnormal{PopInf}}(\widehat{\bm{\omega}})$.
+7:POP-Inf weight $\widehat{\bm{\omega}}$ and the estimator $\widehat{\bm{\theta}}_{\textnormal{POP-Inf}}(\widehat{\bm{\omega}})$.
 3.3 Examples: linear and logistic regression
@@
 yielding a closed-form solution for POP-Inf linear regression
-$\widehat{\theta}_{\textnormal{PopInf}}({\widehat{\bm{\omega}}})$ =$=\left\{\dfrac{1}{n}\sum_{i=1}^{n}{\bf x}_{i}{\bf x}_{i}^{\rm T}+\widehat{\bm{\omega}}\odot\left[\dfrac{1}{N}\sum_{i=n+1}^{n+N}{{\bf x}}_{i}{\bf x}_{i}^{\rm T}-\dfrac{1}{n}\sum_{i=1}^{n}{\bf x}_{i}{\bf x}_{i}^{\rm T}\right]\right\}^{-1}\cdot\left\{\dfrac{1}{n}\sum_{i=1}^{n}{\bf x}_{i}y_{i}+\widehat{\bm{\omega}}\odot\left[\dfrac{1}{N}\sum_{i=n+1}^{n+N}{\bf x}_{i}{\widehat{f}}({\bf z}_{i})-\dfrac{1}{n}\sum_{i=1}^{n}{\bf x}_{i}\widehat{f}({\bf z}_{i})\right]\right\},$ (21)
+$\widehat{\theta}_{\textnormal{POP-Inf}}({\widehat{\bm{\omega}}})$ =$=\left\{\dfrac{1}{n}\sum_{i=1}^{n}{\bf x}_{i}{\bf x}_{i}^{\rm T}+\widehat{\bm{\omega}}\odot\left[\dfrac{1}{N}\sum_{i=n+1}^{n+N}{{\bf x}}_{i}{\bf x}_{i}^{\rm T}-\dfrac{1}{n}\sum_{i=1}^{n}{\bf x}_{i}{\bf x}_{i}^{\rm T}\right]\right\}^{-1}\cdot\left\{\dfrac{1}{n}\sum_{i=1}^{n}{\bf x}_{i}y_{i}+\widehat{\bm{\omega}}\odot\left[\dfrac{1}{N}\sum_{i=n+1}^{n+N}{\bf x}_{i}{\widehat{f}}({\bf z}_{i})-\dfrac{1}{n}\sum_{i=1}^{n}{\bf x}_{i}\widehat{f}({\bf z}_{i})\right]\right\},$ (21)
 where $\widehat{\bm{\omega}}$ is a consistent estimate of optimal $\bm{\omega}_{0}$, which can be obtained by Algorithm 1.
@@
 For logistic regression, $\bm{\psi}(y,{\bf x};{\bm{\theta}})=-{\bf x}y+{\bf x}\gamma_{{\bm{\theta}}}({\bf x})$, where $\gamma_{{\bm{\theta}}}({\bf x})=1/\left\{1+\exp\left(-{\bf x}^{\rm T}{\bm{\theta}}\right)\right\}$.
-Using estimation equation (11), the POP-Inf estimator $\widehat{\bm{\theta}}_{\textnormal{PopInf}}({\widehat{\bm{\omega}}})$ isthe solution to the equation
+Using estimation equation (11), the POP-Inf estimator $\widehat{\bm{\theta}}_{\textnormal{POP-Inf}}({\widehat{\bm{\omega}}})$ isthe solution to the equation
 $\dfrac{1}{n}\sum_{i=1}^{n}\left\{-{\bf x}_{i}y_{i}+{\bf x}_{i}\gamma_{{\bm{\theta}}}({\bf x}_{i})\right\}+\bm{\omega}\odot\left\{\dfrac{1}{N}\sum_{i=n+1}^{n+N}[-{{\bf x}}_{i}\widehat{f}({\bf z}_{i})+{{\bf x}}_{i}\gamma_{{\bm{\theta}}}({\bf x}_{i})]-\dfrac{1}{n}\sum_{i=1}^{n}[-{\bf x}_{i}\widehat{f}({\bf z}_{i})+{\bf x}_{i}\gamma_{{\bm{\theta}}}({\bf x}_{i})]\right\}={\bf 0}.$ (22)
 This does not yield a closed form solution. To obtain an estimator, one can adopt Algorithm 1.
+3.4 Interpretation of the weighting vector
+Next, we provide interpretations of the weighting vector. We build the intuitions using two simple examples, mean estimation and simple linear regression with one-dimensional covariates.
 4 Relationship with Efficient Influence Function
@@
 $\bm{\phi}(y_{i},{\bf x}_{i},{\bm{\theta}})=\dfrac{r}{\pi}{\bf A}^{-1}[\bm{\psi}(y,{\bf x};{\bm{\theta}})-\mathbb{E}\{\bm{\psi}(Y,{\bf x};{\bm{\theta}})\mid{\bf x},{\bf z}\}]+{\bf A}^{-1}\mathbb{E}\{\bm{\psi}(Y,{\bf x};{\bm{\theta}})\mid{\bf x},{\bf z}\}.$ (23)
-Guided by Theorem 1, an EIF-based estimator would be the solution to the empirical equation,
+Proof is contained in Appendix B.2.
+Guided by Proposition 1, an EIF-based estimator would be the solution to the empirical equation,
 ${\bf 0}=\dfrac{1}{n+N}\sum_{i=1}^{n+N}\bm{\phi}(y_{i},{\bf x}_{i},\widehat{\bm{\theta}}_{\textnormal{EIF}}).$ (24)
@@
 Upon simple calculation, we can write the resulting $\widehat{\bm{\theta}}_{\textnormal{EIF}^{*}}$ as the solution to
-$\dfrac{1}{n}\sum_{i=1}^{n}\bm{A}^{-1}\bm{\psi}(y_{i},{\bf x}_{i};{\bm{\theta}})+(1-\pi){\bf A}^{-1}\left\{-\dfrac{1}{n}\sum_{i=1}^{n}\bm{\psi}(\widehat{f}({\bf z}_{i}),{\bf x}_{i};{\bm{\theta}})+\dfrac{1}{N}\sum_{i=n+1}^{n+N}\bm{\psi}(\widehat{f}({\bf z}_{i}),{\bf x}_{i};{\bm{\theta}})\right\}={\bf 0}.$ (26)
+$\dfrac{1}{n}\sum_{i=1}^{n}{\bf A}^{-1}\bm{\psi}(y_{i},{\bf x}_{i};{\bm{\theta}})+(1-\pi){\bf A}^{-1}\left\{-\dfrac{1}{n}\sum_{i=1}^{n}\bm{\psi}(\widehat{f}({\bf z}_{i}),{\bf x}_{i};{\bm{\theta}})+\dfrac{1}{N}\sum_{i=n+1}^{n+N}\bm{\psi}(\widehat{f}({\bf z}_{i}),{\bf x}_{i};{\bm{\theta}})\right\}={\bf 0}.$ (26)
 Similar to the equation (11) that the POP-Inf estimator solves, the $*{}^{*}$-based estimator (26) solves a equation with two parts: one term that can produce a consistent estimator with labeled data only, and the other term utilizes predictions $\widehat{f}$ that is always mean zero.
 If one multiples both sides of (11) with matrix ${\bf A}^{-1}$, then (26) corresponds to that equation weight equals to a vector of $1-\pi$.
-In estimating the outcome mean (Section 2.1) where ${\bf A}$ is essentially one, the $*{}^{*}$-based estimator is the POP-Inf estimator with weight being $1-\pi=N/(n+N)$.
+In estimating the outcome mean (Section 2.1) where ${\bf A}$ is essentially the scalar one, the $*{}^{*}$-based estimator is the POP-Inf estimator with weight being $1-\pi=N/(n+N)$.
 The difference between (11) and (26) clarifies one of the main distinctions between the proposed POP-Inf estimator and an EIF guided estimator.
@@
 Numeric comparisons are provided in Section 6.
+5 Extensions
+Here, we make extensions of POP-Inf to handle more general situations.
+In particular, Section 5.1 describes an inference procedure when the the outcome is fully labeled but covariate ${\bf x}$ is partly labeled and Section 5.2 considers the case where both $y$ and ${\bf x}$ are only partly labeled.
+5.1 ML-predicted covariates
+We adjusted the previously described setup by modifying it to use ML to predict the covariate ${\bf X}$ instead of the outcome $Y$.
+Our data can be divided into two parts: ${\cal L}\cup{\cal U}$.
+The ${\cal L}$ part includes data points $(y_{i},{\bf x}_{i},{\bf z}_{i})$ for $i=1,\cdots,n$ and the ${\cal U}$ part includes data points $(y_{i},{\bf z}_{i})$ for $i=n+1,\cdots,N+n$.
+In addition, we use an external prediction algorithm with the notation $\widehat{q}(\cdot)$.
+This algorithm is applied to ${\bf z}$ to generate the predicted value $\widehat{q}({\bf z})$ of the covariate ${\bf X}$.
+However, it is important to note that $\widehat{q}$ may produce inaccurate or biased predictions of ${\bf X}$.
+The estimand is also defined by the estimating equation (1).
+Similar to our proposal to handle the ML-predicted outcome, we propose our POP-Inf estimator with ML-predicted covariates that solves the equation,
+$\dfrac{1}{n}\sum_{i=1}^{n}\bm{\psi}(y_{i},{\bf x}_{i};{\bm{\theta}})+\widehat{\bm{\omega}}\odot\left\{-\dfrac{1}{n}\sum_{i=1}^{n}\bm{\psi}(y_{i},\widehat{q}({\bf z}_{i});{\bm{\theta}})+\dfrac{1}{N}\sum_{i=n+1}^{n+N}\bm{\psi}(y_{i},\widehat{q}({\bf z}_{i});{\bm{\theta}})\right\}={\bf 0},$ (27)
+where $\widehat{\bm{\omega}}$ is also to minimize the variance of the estimator.
+Such estimation equation is also the estimation equation for the classic estimator with an augmented mean zero term that is indexed by a vector $\bm{\omega}$.
+Therefore, the proposed estimator $\widehat{\bm{\theta}}_{\textnormal{POPInf}}(\bm{\omega})$ is also consistent and asymptotic normally distributed, under mild regularity conditions.
+Corollary 3 states the asymptotic distribution that enables interval estimation with its proof follows from the proof of Theorem 1.
+Corollary 3.
+Under Conditions (C1)-(C3), assuming $\frac{n}{N}\rightarrow\rho$ as $n\to\infty$ and $N\rightarrow\infty$, then the proposed estimator $\widehat{\bm{\theta}}_{\textnormal{POPInf}}(\bm{\omega})$ converges to ${\bm{\theta}}_{0}$ in probability.
+Assuming additionally Condition (C4), we have that as $n\to\infty$ and $N\to\infty$,
+$\sqrt{n}(\widehat{\bm{\theta}}_{\textnormal{POP-Inf}}({\bm{\omega}})-{\bm{\theta}}_{0})\stackrel{{\scriptstyle D}}{{\rightarrow}}\mathcal{N}({\bf 0},{\bm{\Sigma}}^{\prime}(\bm{\omega})).$ (28)
+where ${\bm{\Sigma}}^{\prime}(\bm{\omega})={\bf A}^{-1}{\bf V}^{\prime}(\bm{\omega}){\bf A}^{-1}$, ${\bf V}^{\prime}(\bm{\omega})={\bf M}_{1}+\bm{\omega}\bm{\omega}^{\rm T}\odot({\bf M}_{2}^{\prime}+\rho{\bf M}_{3}^{\prime})-2{\bf I}_{q}\bm{\omega}^{\rm T}\odot{\bf M}_{4}^{\prime}$, ${\bf M}_{2}^{\prime}=\textnormal{Var}_{l}[\bm{\psi}(Y,\widehat{q}({\bf X});{\bm{\theta}}_{0})]$, ${\bf M}_{3}=\textnormal{Var}_{u}[\bm{\psi}(Y,\widehat{q}({\bf X});{\bm{\theta}}_{0})]$, ${\bf M}_{4}^{\prime}=\textnormal{Cov}_{l}[\bm{\psi}(Y,{\bf X};{\bm{\theta}}_{0}),\bm{\psi}(Y,\widehat{q}({\bf X});{\bm{\theta}}_{0})]$.
+5.2 ML-predicted outcome and covariates
+Next, we explore a situation where the outcome and covariates are not directly observed in the unlabeled data.
+Instead, we employ machine learning techniques to predict both.
+Additionally, we possess a much smaller data with both outcome and covariates measured.
+This full data can be categorized into two segments: ${\cal L}\cup{\cal U}$.
+The first segment, ${\cal L}$, includes the data points $(y_{i},{\bf x}_{i},{\bf z}_{i})$ for $i=1,\cdots,n$, while the second segment, ${\cal U}$, comprises the instances of ${\bf z}_{i}$ for $i=n+1,\cdots,N+n$, and $N>>n$.
+The variable ${\bf z}$ is to predict both the outcome and the covariates.
+The ML-predicted outcome and covariates are represented by $\widehat{f}({\bf z})$ and $\widehat{q}({\bf z})$, respectively, where $\widehat{f}$ and $\widehat{q}$ are independently obtained.
+The POP-Inf estimator for this situation is to solve the equation,
+$\dfrac{1}{n}\sum_{i=1}^{n}\bm{\psi}(y_{i},{\bf x}_{i};{\bm{\theta}})+\widehat{\bm{\omega}}\odot\left\{-\dfrac{1}{n}\sum_{i=1}^{n}\bm{\psi}(\widehat{f}({\bf z}_{i}),\widehat{q}({\bf z}_{i});{\bm{\theta}})+\dfrac{1}{N}\sum_{i=n+1}^{n+N}\bm{\psi}(\widehat{f}({\bf z}_{i}),\widehat{q}({\bf z}_{i});{\bm{\theta}})\right\}={\bf 0},$ (29)
+where $\widehat{\bm{\omega}}$ can be obtained by minimizing the variance of the estimator.
+Similarly, we have the consistency and asymptotic normality for the $\widehat{\bm{\theta}}_{\textnormal{POPInf}}(\bm{\omega})$ in Corollary 4 whose proof follows from the proof of Theorem 1.
+Corollary 4.
+Under Conditions (C1)-(C3), assuming $\frac{n}{N}\rightarrow\rho$ as $n\to\infty$ and $N\rightarrow\infty$, then the proposed estimator $\widehat{\bm{\theta}}_{\textnormal{POPInf}}(\bm{\omega})$ converges to ${\bm{\theta}}_{0}$ in probability.
+Assuming additionally Condition (C4), we have that as $n\to\infty$ and $N\to\infty$,
+$\sqrt{n}(\widehat{\bm{\theta}}_{\textnormal{POP-Inf}}({\bm{\omega}})-{\bm{\theta}}_{0})\stackrel{{\scriptstyle D}}{{\rightarrow}}\mathcal{N}({\bf 0},{\bm{\Sigma}}^{\prime\prime}(\bm{\omega})),$ (30)
+where ${\bm{\Sigma}}^{\prime\prime}(\bm{\omega})={\bf A}^{-1}{\bf V}^{\prime\prime}(\bm{\omega}){\bf A}^{-1}$, ${\bf V}^{\prime\prime}(\bm{\omega})={\bf M}_{1}+\bm{\omega}\bm{\omega}^{\rm T}\odot({\bf M}^{\prime\prime}_{2}+\rho{\bf M}^{\prime\prime}_{3})-2{\bf I}_{q}\bm{\omega}^{\rm T}\odot{\bf M}^{\prime\prime}_{4}$, ${\bf M}_{2}^{\prime\prime}=\textnormal{Var}_{l}[\bm{\psi}(\widehat{f}({\bf Z}),\widehat{q}({\bf X});{\bm{\theta}}_{0})]$, ${\bf M}_{3}^{\prime\prime}=\textnormal{Var}_{u}[\bm{\psi}(\widehat{f}({\bf Z}),\widehat{q}({\bf X});{\bm{\theta}}_{0})]$, ${\bf M}_{4}^{\prime\prime}=\textnormal{Cov}_{l}[\bm{\psi}(Y,{\bf X};{\bm{\theta}}_{0}),\bm{\psi}(\widehat{f}({\bf Z}),\widehat{q}({\bf X});{\bm{\theta}}_{0})]$.
 6 Simulations
@@
 Simulation results are based on $1000$ repetitions.
-At each repetition, we calculate standard errors for all estimators based on Theorem 1 and Theorem 1.
+At each repetition, we calculate standard errors for all estimators based on Theorem 1 and Corollary 2.
 6.2 Results
@@
 This suggests that when the prediction performs well, the proposed POP-Inf estimator is still a better choice.
-Similar patterns can be found for $n=500$ (Figure 3).
-In addition, Figure 4 shows boxplots of biases of the POP-Inf estimator across $1000$ repetitions.
+Similar patterns can be found for $n=500$ (Figure 4 in Appendix A).
+In addition, Figure 5 in Appendix A shows boxplots of biases of the POP-Inf estimator across $1000$ repetitions.
 As the sample size increases, the boxes become narrower and hence the variance becomes smaller.
@@
 A ratio lower than one represents a smaller standard error than the classic approach.
-Figure 3: Relative efficiency and coverage rates in simulations with $\gamma\in[0,1]$ and $n=500$.
-In each panel, the x-axis coordinates the data generating parameter $\gamma$ detailed in Section 5.
-Left panels plot the ratios between averaged standard errors of a competing method and the classic approach, with dashed horizontal line intersects $1$ at the y-axis.
-A ratio lower than one represents a smaller standard error than the classic approach.
-Figure 4: Bias of the POP-Inf estimator in simulations for (red) $n=250$ and (purple) $n=500$. Panels are faceted by the estimand and coordinated by $\gamma\in[0,1]$ on the x-axis.
 7 Data Applications
@@
 It would be an interesting future work to carefully discuss relationships between ${\bf X}$ and ${\bf Z}$ and and how they may affect the magnitude of the efficiency improvement.
+The R codes to implement POP-Inf, along with the benchmark methods, and to replicate the simulation and real data analysis, are available at https://github.com/qlu-lab/POPInf.
 Acknowledgments
@@
 We thank the participants of the UW-Madison Statistics Graduate Student Association Student Seminar and the Lu Group Lab Meeting for many useful comments and questions.
+Appendix A Supplementary Figures to Simulations
+Figure 4: Relative efficiency and coverage rates in simulations with $\gamma\in[0,1]$ and $n=500$.
+In each panel, the x-axis coordinates the data generating parameter $\gamma$ detailed in Section 6.
+Left panels plot the ratios between averaged standard errors of a competing method and the classic approach, with dashed horizontal line intersects $1$ at the y-axis.
+A ratio lower than one represents a smaller standard error than the classic approach.
+Figure 5: Bias of the POP-Inf estimator in simulations for (red) $n=250$ and (purple) $n=500$. Panels are faceted by the estimand and coordinated by $\gamma\in[0,1]$ on the x-axis.
+Appendix B Proofs
+Lemma 1.
+Suppose that $X_{1},\ldots,X_{n}$ are i.i.d., and the parameter space $\Theta$ is compact and $L(X,\theta)$ is continuous in $\theta\in\Theta$ almost everywhere.
+Moreover, there exists a function $h(X)$ satisfying $\|L(X,\theta)\|\leq h(X)$ for arbitrary $\theta\in\Theta$ and $\mathbb{E}\{h(X)\}<\infty$, then $\mathbb{E}\{L(X,\theta)\}$ is continuous in $\theta$ and
+$\sup_{\theta\in\Theta}\left\|\frac{1}{n}\sum_{i=1}^{n}L\left(X_{i},\theta\right)-\mathbb{E}\{L(X,\theta)\}\right\|\stackrel{{\scriptstyle P}}{{\rightarrow}}0.$ (32)
+Lemma 1 follows from Tauchen, (1985, Lemma 1) and hence the proof is omitted.
+B.1 Proof of Theorem 1
+For simplicity of notation, we re-write the estimating equation (11) as
+$\Psi(\widehat{\bm{\theta}}_{\textnormal{POP-Inf}})=\Psi_{n}^{Y}+\bm{\omega}\odot(-\Psi_{n}^{\widehat{f}}+\Psi_{N}^{\widehat{f}})={\bf 0},\text{ where}$ (33)
+$\Psi_{n}^{Y}=\dfrac{1}{n}\sum_{i=1}^{n}\psi(y_{i},{\bf x}_{i};\widehat{\bm{\theta}}_{\textnormal{POP-Inf}}),\Psi_{n}^{\widehat{f}}=\dfrac{1}{n}\sum_{i=1}^{n}\psi(\widehat{f}({\bf z}_{i}),{\bf x}_{i};\widehat{\bm{\theta}}_{\textnormal{POP-Inf}}),\text{ and }\Psi_{N}^{\widehat{f}}=\dfrac{1}{N}\sum_{i=n+1}^{n+N}\psi(\widehat{f}({\bf z}_{i}),{\bf x}_{i};\widehat{\bm{\theta}}_{\textnormal{POP-Inf}}).$ (34)
+We first establish consistency following Van der Vaart, (2000, Theorem 5.9) by checking its two conditions.
+The deterministic condition is implied by Condition (C1) and (C2).
+We are left to verify the uniform convergence condition $\|\Psi(\widehat{\bm{\theta}}_{\textnormal{POP-Inf}})-\mathbb{E}(\Psi_{n}^{Y})\|=o_{p}(1)$.
+By Condition (C3) and Lemma 1, we have
+$\|\bm{\omega}\odot\Psi_{n}^{\widehat{f}}-\mathbb{E}(\bm{\omega}\odot\Psi_{n}^{\widehat{f}})\|$ $=o_{p}(1),$ (35)
+$\|\bm{\omega}\odot\Psi_{N}^{\widehat{f}}-\mathbb{E}(\bm{\omega}\odot\Psi_{N}^{\widehat{f}})\|$ $=o_{p}(1),\text{ and}$ (36)
+$\|\Psi_{n}^{Y}-\mathbb{E}(\Psi_{n}^{Y})\|$ $=o_{p}(1).$ (37)
+Hence, by triangular inequality,
+$\|\Psi(\widehat{\bm{\theta}}_{\textnormal{POP-Inf}})-\mathbb{E}(\Psi_{n}^{Y})\|$ (38)
+$\leq$ $\|\Psi(\widehat{\bm{\theta}}_{\textnormal{POP-Inf}})-\Psi_{n}^{Y}\|+\|\Psi_{n}^{Y}-\mathbb{E}(\Psi_{n}^{Y})\|$ (39)
+$=$ $\|-\bm{\omega}\odot\Psi_{n}^{\widehat{f}}+\bm{\omega}\odot\Psi_{N}^{\widehat{f}}\|+\|\Psi_{n}^{Y}-\mathbb{E}(\Psi_{n}^{Y})\|$ (40)
+$\leq$ $\|-\bm{\omega}\cdot\Psi_{n}^{\widehat{f}}+\mathbb{E}(\bm{\omega}\odot\Psi_{n}^{\widehat{f}})\|+\|\bm{\omega}\odot\Psi_{N}^{\widehat{f}}-\mathbb{E}(\bm{\omega}\odot\Psi_{N}^{\widehat{f}})\|+\|\mathbb{E}(\bm{\omega}\odot\Psi_{N}^{\widehat{f}}-\bm{\omega}\odot\Psi_{n}^{\widehat{f}})\|+\|\Psi_{n}^{Y}-\mathbb{E}(\Psi_{n}^{Y})\|$ (41)
+$=$ $o_{p}(1),$ (42)
+which completes the proof of consistency, i.e., $\widehat{\bm{\theta}}_{\textnormal{POP-Inf}}\stackrel{{\scriptstyle P}}{{\rightarrow}}{\bm{\theta}}_{0}$.
+Next, we establish the asymptotic normality.
+Expanding $\Psi(\widehat{\bm{\theta}}_{\textnormal{POP-Inf}})$ at the true value ${\bm{\theta}}_{0}$,
+${\bf 0}=$ $\Psi({\bm{\theta}}_{0})+(\widehat{\bm{\theta}}_{\textnormal{POP-Inf}}-{\bm{\theta}}_{0})\partial\Psi(\bar{{\bm{\theta}}})/\partial{\bm{\theta}}$ (43)
+for some $\bar{{\bm{\theta}}}$ between ${\bm{\theta}}_{0}$ and $\widehat{\bm{\theta}}_{\textnormal{POP-Inf}}$. Multiplying both sides with $\sqrt{n}$ and with a direct calculation,
+$\left\{\partial\Psi(\bar{\bm{\theta}})/\partial{\bm{\theta}}\right\}^{-1}\sqrt{n}(\widehat{\bm{\theta}}_{\textnormal{POP-Inf}}-{\bm{\theta}}_{0})=$ $\sqrt{n}\Psi({\bm{\theta}}_{0})+o_{p}(1).$ (44)
+In order to show the asymptotic normality, by Slutsky’s theorem, it’s sufficient to verify the followings:
+$\|\partial\Psi(\bar{{\bm{\theta}}})/\partial{\bm{\theta}}-\mathbb{E}\{\partial\Psi(\bar{{\bm{\theta}}})/\partial{\bm{\theta}}\}\|=o_{p}(1),$ (45)
+$\sqrt{n}\Psi({\bm{\theta}}_{0})\stackrel{{\scriptstyle D}}{{\rightarrow}}\mathcal{N}({\bf 0},{\bf V}(\omega)).$ (46)
+For (45),
+$\small\|\partial\Psi(\bar{{\bm{\theta}}})/\partial{\bm{\theta}}-\mathbb{E}\{\partial\Psi(\bar{{\bm{\theta}}})/\partial{\bm{\theta}}\}\|\leq$ $\sup_{{\bm{\theta}}\in\Theta}\|\partial\Psi_{n}({\bm{\theta}})/\partial{\bm{\theta}}-\mathbb{E}\{\partial\Psi_{n}({\bm{\theta}})/\partial{\bm{\theta}}\}\|+\|\mathbb{E}\{\partial\Psi_{n}(\bar{{\bm{\theta}}})/\partial{\bm{\theta}}\}-\mathbb{E}\{\partial\Psi_{n}({\bm{\theta}}_{0})/\partial{\bm{\theta}}\}\|.$ (47)
+Noticing $\sup_{{\bm{\theta}}\in\Theta}\|\partial\Psi_{n}({\bm{\theta}})/\partial{\bm{\theta}}-\mathbb{E}\{\partial\Psi_{n}({\bm{\theta}})/\partial{\bm{\theta}}\}\|=o_{p}(1)$ by Condition (C4) and Lemma 1 and $\|\mathbb{E}\{\partial\Psi_{n}(\bar{{\bm{\theta}}})/\partial{\bm{\theta}}\}-\mathbb{E}\{\partial\Psi_{n}({\bm{\theta}}_{0})/\partial{\bm{\theta}}\}\|=o_{p}(1)$ by $\|\bar{{\bm{\theta}}}-{\bm{\theta}}_{0}\|\leq\|\widehat{\bm{\theta}}-{\bm{\theta}}_{0}\|$ and continuous mapping theorem, equation (45) has been verified.
+Meanwhile, (46) holds by Central Limit Theorem.
+The proof of asymptotic normality is completed.
+B.2 Proof of Proposition 1
+The joint log-likelihood for a generic observation $(r,{\bf x},{\bf z},y)$ is
+$l$ $=r{\rm log}p(y\mid{\bf x},{\bf z})+{\rm log}p({\bf x}\mid{\bf z})+{\rm log}p({\bf z})+r{\rm log}(\pi)+(1-r){\rm log}(1-\pi).$ (48)
+Taking the semiparametric approach (Bickel et al.,, 1993; Tsiatis,, 2006), we consider the Hilbert space ${\cal H}$ that contains all one-dimensional zero-mean measurable functions of the observed data with finite variance, equipped with the inner product $\langle h_{1},h_{2}\rangle=\mathbb{E}\{h_{1}(\cdot)h_{2}(\cdot)\}$ where $h_{1},h_{2}\in{\cal H}$.
+To estimate ${\bm{\theta}}$, we regard $p(y\mid{\bf x},{\bf z})$, $p({\bf x}\mid{\bf z})$ and $p({\bf z})$ as nuisance functions.
+Denote their nuisance tangent spaces by ${\cal T}_{y}$, ${\cal T}_{{\bf x}}$, and ${\cal T}_{{\bf z}}$, which are defined as the mean squared closure of the nuisance tangent spaces of parametric submodels spanned by the nuisance score vectors.
+We have that ${\cal T}={\cal T}_{y}\oplus{\cal T}_{{\bf x}}\oplus{\cal T}_{{\bf z}}$ where
+${\cal T}_{y}$ $=[r{\bf a}_{1}(y,{\bf x},{\bf z}):\mathbb{E}\{{\bf a}_{1}(Y,{\bf x},{\bf z})\mid{\bf x},{\bf z}\}={\bf 0}],$ (49)
+${\cal T}_{{\bf x}}$ $=[{\bf a}_{2}({\bf x},{\bf z}):\mathbb{E}\{{\bf a}_{2}({\bf X},{\bf z})\mid{\bf z}\}={\bf 0}],\text{ and}$ (50)
+${\cal T}_{{\bf z}}$ $=[{\bf a}_{3}({\bf z}),\mathbb{E}\{{\bf a}_{3}({\bf Z})\}={\bf 0}],$ (51)
+respectively, and the notation $\oplus$ represents the direct sum of two spaces that are orthogonal to each other.
+We introduce parametric submodels $p_{{\bm{\eta}}_{y}}(y\mid{\bf x},{\bf z})$, $p_{{\bm{\eta}}_{{\bf x}}}({\bf x}\mid{\bf z})$, and $p_{{\bm{\eta}}_{{\bf z}}}({\bf z})$, where ${\bm{\eta}}=[{\bm{\eta}}_{y}^{\rm T},{\bm{\eta}}_{{\bf x}}^{\rm T},{\bm{\eta}}_{{\bf z}}^{\rm T},{\bm{\eta}}_{r}^{\rm T}]^{\rm T}$ is a vector of nuisance parameters.
+Let
+${\bf S}_{{\bm{\eta}}_{y}}(y,{\bf x},{\bf z})=\dfrac{\partial{\rm log}p_{{\bm{\eta}}_{y}}(y,{\bf x},{\bf z})}{\partial{\bm{\eta}}_{y}},\ {\bf S}_{{\bm{\eta}}_{{\bf x}}}({\bf x},{\bf z})=\dfrac{\partial{\rm log}p_{{\bm{\eta}}_{{\bf x}}}({\bf x},{\bf z})}{\partial{\bm{\eta}}_{{\bf x}}},\text{ and }{\bf S}_{{\bm{\eta}}_{{\bf z}}}({\bf z})=\dfrac{\partial{\rm log}p_{{\bm{\eta}}_{{\bf z}}}({\bf z})}{\partial{\bm{\eta}}_{{\bf z}}},$ (52)
+then these scores functions satisfy $r{\bf S}_{{\bm{\eta}}_{y}}(y,{\bf x},{\bf z})\in{\cal T}_{y},\ {\bf S}_{{\bm{\eta}}_{{\bf x}}}({\bf x},{\bf z})\in{\cal T}_{{\bf x}},\text{ and }{\bf S}_{{\bm{\eta}}_{{\bf z}}}({\bf z})\in{\cal T}_{{\bf z}}$.
+Recall the definition of parameter of interest, i.e.,
+[…]
+Taking derivative of ${\bm{\theta}}$ with respect to nuisance parameters,
+$\dfrac{\partial{\bm{\theta}}}{\partial{\bm{\eta}}_{y}^{\rm T}}=$ ${\bf A}^{-1}\dfrac{\partial\mathbb{E}\{\bm{\psi}(Y,{\bf X};{\bm{\theta}})\}}{\partial{\bm{\eta}}_{y}^{\rm T}}={\bf A}^{-1}\mathbb{E}[\bm{\psi}(Y,{\bf X};{\bm{\theta}}){\bf S}_{{\bm{\eta}}_{y}}^{\rm T}(Y,{\bf X},{\bf Z})],$ (54)
+$\dfrac{\partial{\bm{\theta}}}{\partial{\bm{\eta}}_{{\bf x}}^{\rm T}}=$ ${\bf A}^{-1}\dfrac{\partial\mathbb{E}\{\bm{\psi}(Y,{\bf X};{\bm{\theta}})\}}{\partial{\bm{\eta}}_{{\bf x}}^{\rm T}}={\bf A}^{-1}\mathbb{E}\{\bm{\psi}(Y,{\bf X};{\bm{\theta}}){\bf S}_{{\bm{\eta}}_{{\bf x}}}^{\rm T}({\bf X},{\bf Z})\},\text{ and}$ (55)
+$\dfrac{\partial{\bm{\theta}}}{\partial{\bm{\eta}}_{{\bf z}}^{\rm T}}=$ ${\bf A}^{-1}\dfrac{\partial\mathbb{E}\{\bm{\psi}(Y,{\bf X};{\bm{\theta}})\}}{\partial{\bm{\eta}}_{{\bf z}}^{\rm T}}={\bf A}^{-1}\mathbb{E}\{\bm{\psi}(Y,{\bf X};{\bm{\theta}}){\bf S}_{{\bm{\eta}}_{{\bf z}}}^{\rm T}({\bf Z})\}.$ (56)
+Let $\bm{\phi}(y,{\bf x};{\bm{\theta}})=r{\bf a}_{1}(y,{\bf x},{\bf z})+{\bf a}_{2}({\bf x},{\bf z})+{\bf a}_{3}({\bf z})$, where
+${\bf a}_{1}(y,{\bf x},{\bf z})=$ $\dfrac{1}{\pi({\bf z})}{\bf A}^{-1}\left[\bm{\psi}(y,{\bf x};{\bm{\theta}})-\mathbb{E}\{\bm{\psi}(Y,{\bf x};{\bm{\theta}})\mid{\bf x},{\bf z}\}\right],$ (57)
+${\bf a}_{2}({\bf x},{\bf z})=$ ${\bf A}^{-1}\left[\mathbb{E}\{\bm{\psi}(Y,{\bf x};{\bm{\theta}})\mid{\bf x},{\bf z}\}-\mathbb{E}\{\bm{\psi}(Y,{\bf x};{\bm{\theta}})\mid{\bf x},{\bf z}\}\right],\text{ and}$ (58)
+${\bf a}_{3}({\bf z})=$ ${\bf A}^{-1}\mathbb{E}\{\bm{\psi}(Y,{\bf x};{\bm{\theta}})\mid{\bf x},{\bf z}\}.$ (59)
+Then it can be easily verified that $r{\bf a}_{1}(y,{\bf x},{\bf z})\in{\cal T}_{y}$, ${\bf a}_{2}({\bf x},{\bf z})\in{\cal T}_{{\bf x}}$, ${\bf a}_{3}({\bf z})\in{\cal T}_{{\bf x}}$, and
+$\small\mathbb{E}\{\bm{\phi}(Y,{\bf X};{\bm{\theta}})R{\bf S}_{{\bm{\eta}}_{y}}^{\rm T}(Y,{\bf X},{\bf Z})\}=\dfrac{\partial{\bm{\theta}}}{\partial{\bm{\eta}}_{y}^{\rm T}},\ \mathbb{E}\left\{\bm{\phi}(Y,{\bf X};{\bm{\theta}}){\bf S}_{{\bm{\eta}}_{{\bf x}}}^{\rm T}({\bf X},{\bf Z}))\right\}=\dfrac{\partial{\bm{\theta}}}{\partial{\bm{\eta}}_{{\bf x}}^{\rm T}},\text{ and }\mathbb{E}\{\bm{\phi}(Y,{\bf X};{\bm{\theta}}){\bf S}_{{\bm{\eta}}_{{\bf z}}}^{\rm T}({\bf Z})\}=\dfrac{\partial{\bm{\theta}}}{\partial{\bm{\eta}}_{{\bf z}}^{\rm T}}.$ (60)
+Therefore, $\bm{\phi}(y,{\bf x};{\bm{\theta}})$ is the efficient influence function. The proof of Proposition 1 is completed.